### v1

# Simultaneous Modeling of Protein Conformation and Dynamics via Autoregression

Abstract
Understanding protein dynamics is critical for elucidating their biological functions. The increasing availability of molecular dynamics (MD) data enables the training of deep generative models to efficiently explore the conformational space of proteins. However, existing approaches either fail to explicitly capture the temporal dependencies between conformations or do not support direct generation of time-independent samples. To address these limitations, we introduce ConfRover, an autoregressive model that simultaneously learns protein conformation and dynamics from MD trajectories, supporting both time-dependent and time-independent sampling. At the core of our model is a modular architecture comprising: (i) an encoding layer, adapted from protein folding models, that embeds protein-specific information and conformation at each time frame into a latent space; (ii) a temporal module, a sequence model that captures conformational dynamics across frames; and (iii) an SE(3) diffusion model as the structure decoder, generating conformations in continuous space. Experiments on ATLAS, a large-scale protein MD dataset of diverse structures, demonstrate the effectiveness of our model in learning conformational dynamics and supporting a wide range of downstream tasks. ConfRover is the first model to sample both protein conformations and trajectories within a single framework, offering a novel and flexible approach for learning from protein MD data.
00footnotetext: Equal contribution. Work done during their internship at ByteDance. Correspondence to Quanquan Gu <quanquan.gu@bytedance.com>.
1 Introduction
Proteins are flexible molecules that can adopt multiple structures, called conformations. Their ability to transition between different conformations enables biological processes critical to life. Characterizing the behavior of a protein, including its (1) dynamic motions, (2) conformational distribution, and (3) transitions between different states, is crucial for understanding its function and guiding the design of novel proteins [4, 11, 31]. Molecular dynamics (MD) simulations are the “gold standard” for studying protein conformational changes [7, 20, 32]. These simulations use physical models to describe the energy of a protein conformation and the forces acting on its atoms. By simulating atomic motion through classical mechanics and iteratively sampling conformations over time, MD enables researchers to explore different conformations, approximating the conformational distribution at equilibrium, and gain mechanistic insights into protein’s behavior. However, MD simulations are both computationally expensive and technically challenging due to long simulation times and the tendency to become trapped in local energy minima.
Figure 1: Key ideas of ConfRover. (A) Various conformation generation tasks are unified as conditional generation, learned via frame-level autoregressive factorization. Each block represents a frame and conditioning frames are outlined in black. The arrow indicates frame dependency. (B) ConfRover models each frame as a conditional distribution given preceding frames. Sequential dependencies are captured through latent variables $\mathbf{h}$, and conformations are sampled from a diffusion decoder, conditioned on the updated latent.
These challenges have motivated the use of deep generative models to study proteins, leveraging the rich conformational and dynamic information provided by large-scale MD datasets [26, 39]:
(1) Generating the dynamic motions of proteins is a direct analog to MD simulation. Pioneering works modeled this by learning transition probabilities of future conformations from the current state [21, 34, 8]. However, MD trajectories are often non-Markovian due to partially observed coordinates (e.g., protein-only atoms) and environmental coupling (e.g., using Langevin thermostats). To mitigate this, Cheng et al. [6] incorporated higher-order information using multiple context frames, though this requires fix context windows and limits flexibility. Jing et al. [18] instead modeled the joint distribution over the entire trajectory, capturing complex dependencies among frames. Due to training on fixed-length trajectories and the non-autoregressive design, their model has limited inference-time flexibility that cannot generate variable-length trajectories. Li et al. [25] introduced an autoregressive approach for flexible trajectory extension, but its deterministic formulation cannot capture trajectory distributions or generate diverse samples.
(2) Learning the conformational distribution enables sampling time-independent conformations. Several methods [33, 17, 40, 44, 23] train diffusion- or flow-based generative models on conformation ensembles from MD simulation data, bypassing the need for sequential sampling. While effective for generating samples in parallel, they disregard temporal information in MD trajectories and therefore cannot simulate physical motion of proteins.
(3) Conformation interpolation generate transition pathways between different states. Recent works [10, 18] have extended generative modeling to conformation interpolation, where the goal is to generate plausible intermediate samples between known start and end conformational states. Jing et al. [18] framed interpolation as a conditional trajectory generation task, but it requires training task-specific model and has not been evaluated on large proteins.
As these generative problems all stem from the same underlying physical principles and involve sampling from a protein’s conformational space, a natural question arises: can we develop a general framework to learn all of these objectives? We present ConfRover, a framework for simultaneous learning protein conformation distribution and dynamics from MD trajectory data (Figure 1). Our key observation is that, for an MD trajectory $\mathbf{x}^{1:L}$ of length $L$, by adopting a general autoregressive formulation, $p(\mathbf{x}^{1:L})=\prod_{l=1}^{L}p(\mathbf{x}^{l}|\mathbf{x}^{<l})$, where $\mathbf{x}^{<l}$ denotes all preceding frames of $\mathbf{x}^{l}$ in the sequence, we can unify multiple generation objectives as instances of frame generation: (1) generating future frame conditioned on all previous frames, suitable for simulating non-Markovian dynamics; (2) unconditional single-frame generation, $p(\mathbf{x}|\varnothing)$, corresponding to time-independent conformation sampling; (3) flexible frame sequence ordering redefines the dependency structure, enabling tasks such as conformation interpolation.
Our contributions are summarized as follows:
  • We introduce a simple yet general framework to learn both the conformational distribution and dynamics from MD data, supporting multiple generation tasks including trajectory simulation, time-independent conformation sampling, and conformation interpolation.
  • We design a modular architecture that captures temporal dependencies in latent space using efficient causal transformers (i.e., Llama [38]), and directly models conformations in continuous SE(3) space using a diffusion decoder, avoiding discretization structure into tokens.
  • Experiments show strong capabilities of ConfRover: it outperforms MDGen [18] in trajectory simulation, matches the performance of AlphaFlow [17] and ConfDiff [40] in time-dependent generation, and can effectively sample conformations interpolating two endpoints.
2 Background
2.1 Data Generation from Molecular Dynamics
Molecular dynamics describes the motion of molecules through Newtonian mechanics $M\ddot{\mathbf{x}}=-\nabla U(\mathbf{x})$, where $\mathbf{x}$ denotes coordinates of atoms in the system, $M$ is the atomic mass, $U(\mathbf{x})$ is the potential energy of the configuration and $-\nabla U(\mathbf{x})$ represents the forces acting on atoms. In practice, stochastic and frictional forces are integrated to model energy exchange with the environment and maintain temperature control of the system, converting the equations of motion to a Langevin process:
$M\ddot{\mathbf{x}}=-\nabla U(\mathbf{x})-\gamma M\dot{\mathbf{x}}+\sqrt{2M\gamma k_{B}T}\boldsymbol{\eta}(t),$
where $\gamma$ is the friction coefficient, $k_{B}$ is the Boltzmann constant, $T$ is the temperature, and $\boldsymbol{\eta}(t)$ is a Gaussian noise term delta-correlated in time $\langle\eta_{i}(t)\eta_{j}(t^{\prime})\rangle=\delta_{ij}\delta(t-t^{\prime})$. Sampling from this stochastic process generates a time evolution of system configurations. Over time, the samples converges to the Boltzmann distribution $p(\mathbf{x})\propto\exp(-U(\mathbf{x})/k_{B}T)$. The trajectory of protein coordinates, $\mathbf{x}^{1:L}_{\text{prot}}=(\mathbf{x}_{\text{prot}}^{1},\mathbf{x}_{\text{prot}}^{2},\dots,\mathbf{x}_{\text{prot}}^{L})$, is extracted and saved at prescribed simulation intervals, providing both distributional and kinetic information the protein conformational dynamics. For simplicity, we omit the subscript ‘prot’ and use $\mathbf{x}$ to denote protein coordinates throughout the paper.
2.2 Protein Representations
[…]
2.3 SE(3)-Diffusion Models for Protein Conformation Generation
Diffusion generative models are capable of learning complex data distributions. Training involves progressively corrupting data with noise and learning to reverse this process through denoising, thereby modeling the original data distribution [13, 35]. Recently, diffusion models operating in SE(3) space have been proposed to model protein backbone structures [43, 40]. Below, we provide a brief overview of diffusion model and defer the detailed SE(3) formulation to Appendix B:
Given a protein backbone conformation as $\mathbf{x}_{0}=(\mathbf{T}_{0},\mathbf{R}_{0})\in\mathrm{SE(3)}^{N}$, and conditioned on the protein identity (omitted in the equations for clarity), we aim to train a neural network to jointly estimate the score functions of the reverse-time marginal distributions at varying diffusion time $t$, $s_{\theta}(\mathbf{x}_{t},t)\approx\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{x}_{t})$. This model is trained using the denoising score matching (DSM) loss:
$\mathcal{L}_{\text{DSM}}=\mathbb{E}_{\mathbf{x}_{0},\mathbf{x}_{t},t}\left[\lambda(t)\|s_{\theta}(\mathbf{x}_{t},t)-\nabla_{\mathbf{x}_{t}}\log p_{t|0}(\mathbf{x}_{t}|\mathbf{x}_{0})\|^{2}\right].$
Here $p_{t|0}(\mathbf{x}_{t}|\mathbf{x}_{0})$ is the forward transition kernel defined in the SE(3) space, $\mathbf{x}_{t}=(\mathbf{T}_{t},\mathbf{R}_{t})$ is the noisy data at time $t$, and $\lambda(t)$ is a time-dependent weight. During inference, DPM generates clean conformations from random noise by simulating the reverse diffusion process with the learned score network $s_{\theta}(\mathbf{x}_{t},t)$.
3 ConfRover
3.1 Modeling MD Trajectories through Autoregression
Autoregressive generative models factorize the distribution of a sequence as a series of conditional generations over frames. We cast this idea to MD trajectories, modeling a sequence of $L$ frames as:
$p(\mathbf{x}^{1:L}|\mathcal{P})=\prod_{l=1}^{L}p(\mathbf{x}^{l}|\mathbf{x}^{<l},\mathcal{P}),$ (1)
where $\mathbf{x}^{<l}$ is the preceding frames and $\mathcal{P}$ denotes the protein-specific conditioning input.
Despite its simplicity, this formulation naturally supports multiple learning objectives in protein conformation modeling. In its base form, it models temporal dependencies across the frames, learning to generate the trajectory. When $L=1$, it removes the frame context and reduces to a single-frame distribution $p(\mathbf{x}|\mathcal{P})$, learning to direct sample time-independent conformations. In addition, the sequential dependency in Equation (1) can be extended to any desired frame-conditioned generation tasks. By prepending conditioning frames $\mathcal{K}$ to the sequence, we train the model to learn any conditional generation $p(\mathbf{x}^{1:L}|\mathcal{K},\mathcal{P})$, including conformation interpolation by setting $\mathcal{K}=\{\mathbf{x}^{1},\mathbf{x}^{L}\}$. Similar idea was applied in text infilling tasks by shuffling the order of text contexts [3].
After defining the main learning objectives for trajectory simulation, single-frame (time-independent), and conformation interpolation, we describe how to effectively model autoregressive dependencies over protein conformations in Section 3.2. More critically, in Section 3.3, we explain how to adapt sequence models, traditionally designed for discrete tokens, to the continuous space of protein conformations. Lastly, we introduce a specific choice of architectures of ConfRover in Section 3.4.
3.2 Latent Causal Modeling
We propose a modular design composed of an encoder, a latent sequence model, and a stochastic decoder. This design enables the use of modern causal transformers, such as in Llama [38], to efficiently capture sequential dependencies between frames in the latent space (Figure 2). During training, the input sequence is shifted by one frame and a mask token “[M]” is prepended; the generation process also begins with the mask token and conditioning frames (e.g., $\mathbf{x}^{1}$).
Figure 2: Causal sequence model to generate trajectory ($\hat{\mathbf{x}}^{2},\dots$) from the mask token “[M]” and the conditioning frame $\mathbf{x}^{1}$. Each frame only attend to its previous frames. Attention activations for $\hat{\mathbf{x}}^{3}$ are highlighted in orange.
To model $p(\mathbf{x}^{l}|\mathbf{x}^{<l},\mathcal{P})$, the context frames $\mathbf{x}^{<l}$ are first encoded into intermediate latent states $\mathbf{h}^{<l}=(\mathbf{h}^{1},\dots,\mathbf{h}^{l-1})$ using a shared encoder network with protein-specific condition $\mathcal{P}$:
$\mathbf{h}^{i}$ $=f_{\eta}^{\text{enc}}(\mathbf{x}^{i},\mathcal{P}),\quad i=1,2,\dots,l-1.$ (2)
A temporal module is then used to capture the sequential dependencies between the frame latents. Causal attention is used to ensure the latent for frame $\mathbf{h}^{l}$ is only updated by its preceding frames:
$\mathbf{h}^{l}_{\text{updated}}$ $=f_{\xi}^{\text{temp}}\left(\mathbf{h}^{1},\mathbf{h}^{2},\dots,\mathbf{h}^{l-1}\right).$ (3)
Since that both the encoder and the temporal module are chosen to be deterministic, $p(\mathbf{x}^{l}|\mathbf{x}^{<l})$ reduces to a conditional generation over the updated latent, realized through a probabilistic decoder:
$p(\mathbf{x}^{l}|\mathbf{x}^{<l},\mathcal{P})$ $=p_{\theta}^{\text{dec}}(\mathbf{x}^{l}|\mathbf{h}^{l}_{\text{updated}}).$ (4)
Details on this latent modeling is provided in Appendix C. As a result, we train a model $(\eta,\xi,\theta)$ by jointly optimizing the three modules $f_{\eta}^{\text{enc}}$, $f_{\xi}^{\text{temp}}$, and $p_{\theta}^{\text{dec}}$.
This setup easily accommodates learning single-frame distribution: by replacing all input frames with a mask token and using identity attentions, where each frame only attends to itself, we effectively disable inter-frame information flow. This trains the model to directly sample conformations. A similar strategy has been used in image-video training [27, 15].
3.3 Training Autoregressive Model with SE(3) Diffusion Loss
The main challenge in applying autoregressive modeling to conformation trajectories lies in representing the continuous distribution of protein conformations within a framework typically used for discrete token sequences. While some studies have approached this by discretizing the protein structural space into discrete tokens [12, 28, 29], such approaches inherently suffer from discretization error, which can lead to suboptimal performance in modeling protein conformations.
Instead, we propose to directly model the continuous conformational space using diffusion probabilistic models and employ the DSM loss for autoregressive model training, similar to Li et al. [24]. Specifically, we perform DSM loss training in SE(3) space.
Given a clean frame $\mathbf{x}^{l}_{0}=(\mathbf{T}_{0}^{l},\mathbf{R}_{0}^{l})$, its latent embedding with temporal context $\mathbf{h}^{l}$ (omit subscript “update” for clarity), the forward transition kernels $p_{t|0}(\mathbf{T}_{t}^{l}|\mathbf{T}_{0}^{l})$ and $p_{t|0}(\mathbf{R}_{t}^{l}|\mathbf{R}_{0}^{l})$ for the translation and rotation component of SE(3), and a score network to jointly estimate the translation and rotation scores $s_{\theta}(\mathbf{T}_{t}^{l},\mathbf{h}^{l},t)$ and $s_{\theta}^{r}(\mathbf{R}_{t}^{l},\mathbf{h}^{l},t)$, the loss is defined as
$\mathcal{L}_{\text{DSM}}^{\text{SE(3)}}$ $=\mathbb{E}\left[\lambda(t)\|s_{\theta}(\mathbf{T}_{t}^{l},\mathbf{h}^{l},t)-\nabla_{\mathbf{T}_{t}^{l}}\log p_{t|0}(\mathbf{T}_{t}^{l}|\mathbf{T}_{0}^{l})\|^{2}\right]$
$\quad\quad+\mathbb{E}\left[\lambda^{r}(t)\|s_{\theta}^{r}(\mathbf{R}_{t}^{l},\mathbf{h}^{l},t)-\nabla_{\mathbf{R}_{t}^{l}}\log p_{t|0}(\mathbf{R}_{t}^{l}|\mathbf{R}_{0}^{l})\|^{2}\right],$ (5)
where the expectation is taken over the diffusion time $t$ and noisy structure $\mathbf{x}_{t}^{l}=(\mathbf{T}_{t}^{l},\mathbf{R}_{t}^{l})$ sampled from the forward process. Gradients with respect to $\mathbf{h}^{l}$ are then backpropagated to update the weights in the temporal module $f_{\xi}^{\text{temp}}$ and encoder $f_{\eta}^{\text{enc}}$. During inference, we decode each frame autoregressively by performing reverse sampling as in Equation (6), replacing the scores with estimated values from $s_{\theta}(\mathbf{T}_{t}^{l},\mathbf{h}^{l},t)$ and $s_{\theta}^{r}(\mathbf{R}_{t}^{l},\mathbf{h}^{l},t)$.
3.4 Model Architecture
An overview of model architecture is shown in Figure 3, with detailed illustrations of each module provided in Appendix D.1.
Figure 3: Architecture overview. (A) Encoding Layer embeds protein sequence and input structure to each frame as a frame latent representation $\mathbf{h}^{l}$, comprised of single and pair embeddings; (B) The Trajectory Module then updates frame latent $\mathbf{h}^{l}$ using interleaved structural and temporal update blocks; (C) A diffusion-based Structure Decoder learns to denoise a noisy conformation conditioned on the updated frame latent $\mathbf{h}^{l}$; during inference, it samples conformations from the prior distribution. See Appendix D.1 for details.
Encoding Layer. A FoldingModule, parameterized by a pretrained OpenFold model [2], extracts protein-specific embeddings $\mathcal{P}$ consisting of a single representation ($\mathbf{s}$) and a pair representation ($\mathbf{z}$), shared across frames. For each frame, a FrameEncoder, adapted from the template module used in prior works [19, 17], encodes pairwise distance of pseudo–Cβ atoms via triangular updates and merges this frame pair representations $\mathbf{z}_{\text{Frame}}^{l}$ with the protein pair representation $\mathbf{z}$. The resulting frame latent embedding, $\mathbf{h}^{l}=[\mathbf{s}^{l},\mathbf{z}^{l}]$, are invariant to global translation and rotation of the conformations, and are passed to the Trajectory Module. Following causal sequence modeling, a masked frame token “[M]” is introduced by zeroing out the pseudo-Cβ pairwise distances to remove structural information.
Trajectory Module. The Trajectory Module models structural and temporal dependencies across frames, updateing each frame’s embedding based on its preceding frames. It consists of interleaved StructualUpdate and TemporalUpdate layers that operate on the frame-wise latent embeddings. StructralUpdate incorporates Pairformer layers, a core architecture in protein structure modeling, to update single and pair embeddings through triangular operations [2]. TemporalUpdate employs a Llama-based causal transformer layer for channel-wise self-attention over the sequence of frame embeddings. Each channel in the single and pair embeddings is updated independently. Frame indices are encoded using Rotary Position Embedding [37]. This interleaved design enables efficient updates while maintaining flexibility in modeling sequential dependencies.
Structure Decoder. The updated latent embeddings from Trajectory Module serve as conditioning signals for generating the conformation at each frame. For the SE(3) diffusion model described in Section 2.3, we adopt ConfDiff [40] as the DiffusionDecoder to generate 3D conformations. ConfDiff composes of layers of Invariant Point Attention and Transformer (on single embeddings) to collectively update the residue SE(3) rigids, as well as single and pair embeddings. Trained with denoising score matching in Equation (5), the DiffusionDecoder learns to iteratively denoise noisy frame structures drawn from a prior SE(3) distribution, conditioned on the frame latent embeddings, to generate accurate backbone conformations of the frame. To reconstruct full-atom geometry, we additionally predict the 7 torsional angles $(\phi,\psi,\omega,\chi_{1},\dots,\chi_{4})$ using a light-weight AngleResNet, for the coordinates of backbone oxygen atom and side-chain atoms.
4 Experiments
Dataset. We evaluate model performance on ATLAS [39], a large-scale protein MD dataset covering $\sim$1300 proteins with diverse sizes and structures. For each protein, it contains triplicated 100 ns simulation trajectories. All models are trained on training trajectories and evaluated on test trajectories split by protein identity [17, 40, 18]. This presents a challenging task for assessing the generalization to unseen protein structures and dynamics.
Model training. All modules are trained from scratch, except for the FoldingModule (frozen OpenFold weights) and the DiffusionDecoder (initialized from ConfDiff). During training, sub-trajectories of length $L=8$ with varying strides ($1\sim 1024$ MD snapshots saved at 10 ps intervals) are sampled to enable learning across multiple timescales. For the main ConfRover model, we adopt a hybrid training strategy using a 1:1 ratio of trajectory and single-frame training objectives; For conformation interpolation task, we continue training the model with a 1:1:1 ratio of trajectory, single-frame and interpolation objectives. See Appendix D.2 for training details.
Baselines. We compare ConfRover with state-of-the-art deep learning models for each task: For trajectory simulation, we compare against MDGen, a flow-based non-autoregressive trajectory model trained on ATLAS; For time-independent generation, we evaluate against AlphaFlow and ConfDiff, flow- and diffusion-based conformation generation models finetuned on ATLAS; For conformation interpolation, no existing baseline is available for large proteins, so we focus on analyzing results of our model. As this is an emerging research area, baseline models remain limited. While recent works on trajectory simulation [21, 8, 6] and conformation interpolation [10, 18] have been proposed, they are not tailored for large proteins as in ATLAS or are not publicly available at the time of manuscript preparation. See Appendix A for more details.
4.1 Trajectory Simulation
Since trajectories from both MD and models are stochastic samples, directly comparing them using frame-wise error, such as root-mean-square-deviation (RMSD) between atomic coordinates, is not appropriate. Therefore, we evaluate the model’s ability to recover trajectory dynamics from two perspectives: (1) how well it captures the magnitude of conformation changes across varying start conformations and timescales; (2) how well it recovers the conformational states and principal dynamic modes observed in long-time MD simulations.
Table 1: Pearson correlations of conformation changes between sampled and reference trajectories in multi-start. The better scores are highlighted in bold.
[…]
Traj. Frame $\Delta$Frame
MDGen 0.55 0.45 0.40
ConfRover 0.77 0.63 0.53
PCA 2D
Traj. Frame $\Delta$Frame
MDGen 0.16 0.11 0.10
ConfRover 0.75 0.50 0.43
[…]
ConfRover shows superior performance in recovering the magnitude of conformational changes. We report the Pearson correlation of measured conformation changes between model-generated and reference trajectories in Table 1, with additional results in Appendix E.1. Compare with MDGen, ConfRover shows a significant improvement in correlation scores, mean absolute error and structural quality (Table 8), indicating its stronger ability to recover the magnitude of conformation changes across different starting conditions in the conformational space. The greater difference observed in PCA highlights that ConfRover more accurately captures conformational changes along the feature dimensions most relevant to the structural variance observed in MD. Figure 4 visualizes ensembles of conformations in generated trajectories, with additional examples in Figure 12. ConfRover exhibit more notable conformation changes than MDGen, and reflects the major movements in the structured and loop domains observed in the MD reference.
MDGen is a non-autoregressive model trained on trajectories of length $L=250$. Adjusting its inference setup results in degraded conformation. To ensure fair comparison, we use the original inference setting ($S=40,L=250$) and downsampled the trajectories for evaluation. To confirm that this post-processing step does not introduce artifacts, we also trained MDGen models under the evaluation setups. The results are consistent with the downsampled version (see Appendix E.5).
Figure 4: Visualization of six proteins from multi-start. Trajectory conformations are colored by their secondary structures and superposed to show the dynamic ensemble. MDGen primarily exhibits local movements, whereas ConfRover captures conformations changes similar to MD simulations.
Table 2: State recovery in 100 ns simulation. Mean values over the first two PCA components are reported. Precision scores for all models are close to 1 and omitted. The better scores are highlighted in bold. MD 100ns is the oracle and excluded in the comparison.
JSD ($\downarrow$) Recall ($\uparrow$) F1 ($\uparrow$)
MD 100ns 0.31 0.67 0.79
MDGen 0.56 0.30 0.44
ConfRover 0.51 0.42 0.58
Assessing long trajectory generation on 100 ns simulation. We further evaluate model’s ability to recover conformational states and principal dynamics of proteins. For each of protein, we simulate a trajectory of $L=80$ frames at stride $S=120$, approximating the 100 ns MD simulation in ATLAS. To assess state recovery, model-generated conformations are projected into a reduced PCA space and compared with the reference trajectory. Specifically, we discretize each principal component into 10 evenly sized “states” and measure the distribution similarity using Jensen-Shannon Distance (JSD). We also compute precision, recall, and F1-score on whether sample conformations fall within these known states [30, 40, 44]. To evaluate dynamic mode recovery, we perform time-lagged independent component analysis (tICA) at varying lag times on both reference and sample trajectories. We then compute Pearson correlations between the per-residue contribution to the leading components, based on the tICA coefficients. See Appendix E.2 for details. We include one of the triplicate MD trajectories—excluded from ground-truth evaluation—as an “oracle” reference, denoted as MD 100ns, representing the performance expected if the model were as accurate as an MD simulation run.
ConfRover recovers more conformational states than MDGen and accurately captures the principal dynamics. As shown in Table 4.1, ConfRover outperforms MDGen in state recovery, achieving lower JSD, higher recall and F1 scores, showing its improved ability to capture diverse conformations. Additionally, ConfRover shows clear advantage in capturing the principal dynamic modes across varying lag times, performing even comparably to the MD oracle (Figure 5A). This results suggest ConfRover can learn and generalize dynamics to unseen proteins and still capture the most important dynamic modes. We visualize simulated trajectories in the PCA space in Figure 5B and Figure 13. These examples again confirm that ConfRover is more capable of sampling over the conformational space of the proteins and covering diverse conformations. Yet, we also observe some cases where MD 100ns overcame the energy barrier and achieved more remote states while ConfRover did not (e.g., 7NMQ-A in Figure 5B).
Summary. These experiments demonstrate that ConfRover outperforms the current state-of-the-art model in trajectory simulation, effectively learning protein dynamics from MD data and generalizing well to unseen proteins. While a gap remains compared to the oracle MD 100ns, particularly in state recovery, the improvement narrows the gap between deep generative models and established simulation methods.
Figure 5: Results from 100 ns simulation. (A) Correlations of principal dynamic modes between sample and reference trajectories, evaluated at varying lag time. (B) Examples trajectories illustrating the states explored by different methods (downsampled by 5 frames for visualization). The blue background indicates the density of the ground-truth conformation distribution from MD reference.
4.2 Time-independent Conformation Sampling
We evaluate the time-independent sampling performance of ConfRover, following the benchmark setup in Ye et al. [42]. For each protein, 250 independent conformations are generated and compared to MD reference ensembles. Key metrics are summarized in Table 3 with full results in Appendix E.3.
ConfRover matches the performance of state-of-the-art specialized models. Compared to AlphaFlow and ConfDiff, both specialized for time-independent conformation sampling, ConfRover achieves overall comparable performance and outperforms at least one of the SOTA models in five evaluation criteria. This demonstrate that ConfRover, despite being a general-purpose model capable of trajecotry generation, also performs strongly in sampling independent conformations that approximate the equilibrium distribution from MD simulation. In contrast, MDGen, which is trained solely for trajectory generation, shows suboptimal results with sequentially sampled conformations.
Table 3: Time-independent generation benchmark. AlphaFlow and ConfDiff are specialized SOTA models. The best scores are highlighted in bold and the second-best scores are underlined. *MDGen does not support time-independent sampling; its sequential sampling result are evaluated.
Pairwise
RMSD
$r$($\uparrow$)
Per target
RMSF
$r$($\uparrow$)
RMWD
($\downarrow$)
MD
PCA
$\mathcal{W}_{2}$($\downarrow$)
Joint
PCA
$\mathcal{W}_{2}$($\downarrow$)
Weak
contacts
$J$ ($\uparrow$)
Transient
contacts
$J$($\uparrow$)
Exposed
residue
$J$($\uparrow$)
AlphaFlow 0.53 0.85 2.64 1.55 2.29 0.62 0.41 0.69
ConfDiff 0.59 0.85 2.75 1.41 2.27 0.63 0.39 0.65
MDGen* 0.41 0.74 2.81 1.95 2.38 0.50 0.28 0.57
ConfRover-Traj 0.47 0.85 2.83 1.41 2.30 0.53 0.36 0.60
ConfRover 0.51 0.85 2.62 1.39 2.28 0.61 0.38 0.67
Effect of hybrid training. Without explicit single-frame training, the model primarily learns time-dependent generation, with only the first frame of each trajectory learning to generate conformation unconditionally (i.e., from a masked token input). To test the importance of hybrid training, we ablated the single-frame objective and trained a variant, ConfRover-traj, solely on trajectory generation. As shown in Table 3, while this variant still outperforms time-dependent results from MDGen, it shows decreased performance across several metrics compared to ConfRover. This highlights the importance of hybrid training in balancing the learning objectives and enhancing the model’s for generating independent conformations.
4.3 Conformation Interpolation
[…]
Training on the interpolation objective enables smooth interpolation between conformations. As shown in Figure 6A, the distance to the start frame increases while the distance to the end frame decreases with frame index, indicating smooth and directed transitions. Without explicit interpolation training, the original ConfRover (dashed lines in Figure 6A) generates trajectory that do not progress towards the end state. Figure 6B visualizes intermediate structures and transition pathways in PCA space, showing that intermediate conformations from ConfRover-interp closely resemble those in the MD reference. In contrast, as shown in Figure 15, the original ConfRover can miss key transitions and fails to reach the end state. Additional results and visualizations are provided in Appendix E.4. These results highlight the effectiveness of our interpolation training strategy: by adjusting the dependency order in the sequence model, ConfRover-interp learns to generate smooth transitions between two conformations.
[…]
4.4 Extension to Masked Sequence Modeling
Figure 7: Autoregressive generation with masked language modeling. $\mathbf{x}^{1}$ is the initial frame for simulation. Masked input frames are iteratively replaced by the generated frames. Attention activations for the current predicting frame $\hat{\mathbf{x}}^{3}$ are highlighted in orange.
While the causal transformers used in ConfRover enables efficient autoregressive training and generation, proposed framework can be extended to other sequence modeling paradigm, such as masked sequence modeling. Similar to Li et al. [24], sequence models trained with bidirectional attention and masked sequence modeling can perform autoregressive generation by iteratively predicting frames in a specified order. At each iteration, the generated frame replaces its corresponding mask and joins the input sequence for the next iteration. We trained a variant ConfRover-mask following this approach. Specifically, we replaced causal attention with bidirectional attention in the transformer and implemented a scheduled masking strategy for training, linearly increasing the mask rate from 45% to 88%.
ConfRover-Mask learns conformation changes in trajectories but lacks inference flexibility. We evaluated ConfRover-mask on the multi-start benchmark. As shown in Table 4, ConfRover-mask outperforms the baseline, though it performs slightly worse than the ConfRover (causal) model. These results demonstrate the flexibility of our framework to extend to a broader range of sequence modeling approaches. However, masked sequence modeling presents several practical limitations. Similar to non-autoregressive model, ConfRover-mask lacks the flexibility to generate trajectories of variable lengths. We observed structural degradation when inferring with lengths differ significantly from the training setup, such as setting $L=1$ for time-independent conformation sampling. As a result, this approach is not well-suited for generating long trajectories without resorting to workarounds such as sliding window prediction with window sizes matching the training window size. While training improvements, such as varying training window sizes, may help address these issues, we leave such exploration to future work.
Overall, the comparison between ConfRover (causal) and ConfRover-mask highlights the versatility of causal sequence modeling, which can be more readily adapted to diverse generation scenarios.
Table 4: Results of ConfRover-mask on the Pearson correlations of conformation changes between sampled and reference trajectories in multi-start. The better scores are highlighted in bold.
[…]
Traj. Frame $\Delta$Frame
MDGen 0.55 0.45 0.40
ConfRover 0.77 0.63 0.53
ConfRover-mask 0.71 0.60 0.49
PCA 2D
Traj. Frame $\Delta$Frame
MDGen 0.16 0.11 0.10
ConfRover 0.75 0.50 0.43
ConfRover-mask 0.73 0.49 0.41
5 Conclusions and Limitations
We introduce ConfRover, a general framework for learning protein conformational dynamics from MD trajectory data. Through autoregressive factorization, ConfRover supports three tasks in a unified manner: trajectory simulation, time-independent sampling, and conformation interpolation. This formulation reflects the temporal nature of MD while naturally encompassing conditional and unconditional frame-level generation. Extensive experiments and analyses highlight several empirical advantages: (1) ConfRover outperforms the current state-of-the-art in trajectory simulation, accurately capturing dynamic magnitude, state recovery, and principal motions; (2) Despite being a multi-purpose model, it achieves competitive performance in time-independent sampling compared to specialized methods; (3) With simple sequence reordering, ConfRover effectively learns to interpolate between conformations.
Nevertheless, several limitations still remain: (1) Trajectory simulation and interpolation are emerging tasks with limited baselines. We hope this work encourages further development and open-sourcing of models for comprehensive benchmarking; (2) Some evaluation metrics are still preliminary. While interpolation results are intuitive, they would benefit from more meaningful benchmarks reflecting functional state transitions. Evaluation is also constrained by the ATLAS dataset, which is limited to 100 ns simulations and mainly local dynamics. (3) Although ConfRover narrows the gap with classical MD, it still falls short in fully capturing the conformational space. Future gains may come from scaling training data, using more efficient architectures, and leveraging additional information from MD, such as energy information. (4) Finally, while the triangular updates in the structural modules ensure high accuracy, their computational cost limits scalability to larger proteins and longer trajectories. Despite these challenges, ConfRover demonstrates the promise of autoregressive models in molecular simulation, offering a unified, efficient, and extensible approach to modeling protein dynamics.
6 Acknowledgments
We thank Zaixiang Zheng for the technical discussions and valuable feedback during the preparation of this manuscript.
Appendix A Extended Discussions
A.1 Related Work
Deep Generative Models for MD Trajectories.
Recent works have explored protein trajectory generation as a surrogate for MD simulations. Models such as Timewarp [21], ITO [34], and EquiJump [8] learn stochastic transport functions to sample future conformations at a lagged time (longer than MD intervals), reducing the computational costs of long-timescale simulations. However, these methods assume Markovian dynamics by relying solely on the current state for prediction, which may not be suitable for non-Markovian dynamics common in protein MD data. To capture higher-order dependencies between the frames, AlphaFolding [6] incorporates history frames via “motion nodes”, but it requires a fixed context window. MDGen [18] takes a different approach by directly modeling the joint distributions of frames in a trajectory and learning frame dependencies through “masked frame modeling”, similar to masked language modeling. However, its key-frame parameterization requires separate models for different tasks, and its non-autoregressive paradigm limits flexible generation (e.g., not compatible for generating trajectories with variable lengths). GST [25] applies autoregression for future frame prediction, enabling variable-length conditioning context and prediction horizons. While the autoregressive approach is conceptually similar to ConfRover, their work differs in several key aspects: it performs deterministic prediction rather than generative trajectory sampling; it employs a graph-based architecture with fixed structural priors from an adjacency graph, instead of full attention across all residues; it is trained and evaluated on a single protein instead of diverse proteins from ATLAS under a transferable setting.
Beyond forward trajectory simulation, generative models have also been applied to conformation interpolation, that is, generating the intermediate trajectories between two conditioned states. The non-autoregressive framework of MDGen [18] can be extended to sample transition pathways between such states; however, its key-frame parameterization requires training a separate model for this task, and it has not been tested on large proteins like those in ATLAS. Du et al. [10] proposed a simulation-free objective for transition-pathway sampling based on Doob’s $h$-transform, but their approach has only been validated on numerical models and the small protein Chignolin. Its generalizability to larger, more diverse proteins remains unassessed.
Notably, the above models focus on learning temporal dependencies between frames and do not support direct, time-independent conformation sampling from the learned distribution. In contrast, ConfRover is a general framework that learns both the trajectory generation tasks as well as direct sampling of independent protein conformations.
Deep Learning Models for Molecular Conformations.
Another line of work focuses on direct sampling of conformations in a time-independent manner. Early efforts include perturbing the input to folding models (e.g., AlphaFold) [9, 41, 36] or perturbing the conformation using a structural diffusion model [30]. However, these models are trained solely on static PDB structures and do not explicitly model the conformational distribution. Recent works have shifted to deep generative paradigms that directly learn protein-specific conformational distributions [16, 44, 17, 40, 29, 23]. Several models in this category, including AlphaFlow, ConfDiff, and BioEmulator, fine-tune pretrained structure models on large-scale MD datasets, enhancing their ability to capture conformational distributions. A related approach trains normalizing flow models to approximate the Boltzmann distribution [33, 22], but their invertibility constraints limit scalability and transferability beyond small molecules and peptides. While these methods can generate time-independent conformations, they overlook temporal relationships and do not capture the kinetic aspects of protein dynamics. ConfRover handles both time-dependent trajectory generation and time-independent conformation learning.
Image-Video Generation.
The challenge of modeling protein dynamics conceptually parallels tasks in image and video generation, requiring both data distribution learning and temporal modeling. Recent advances in video generation offer valuable insights into addressing these challenges. Given limited video data, extending image generative models to video has proven effective. Several works [15, 5, 14] achieve this by incorporating temporal attention layers, enabling frame-to-frame communication. Disabling temporal attention reverts them to image models, allowing flexible training across both modalities. These approaches efficiently model time correlations without explicitly tracking offsets between frames. Meanwhile, the extension of autoregressive language models to image and video domains has shown strong potential for sequential generation in different data modalities. Li et al. [24] integrates language models’ sequential modeling with diffusion models’ ability to model continuous distributions, showing that discrete tokens are not essential for autoregressive models. MarDini [27] extends the concept to video generation with efficient llama-style temporal planning and high-resolution video generation via a diffusion decoder. By applying masked “frame” modeling, MarDini allows the model to learn flexible temporal relationships and enables diverse tasks such as frame interpolation. ConfRover differs from these works from video models in several aspects: it employs SE(3) diffusion for 3D structure generation; by using an autoregressive paradigm, it explicitly decouples the diffusion generation from temporal modeling, unlike the spatiotemporal denoising process in MarDini; in addition, the causal autoregressive framework enables more flexible trajectory generation with variable lengths.
A.2 Training and Inference Cost
ConfRover models in this work contains 19.6 M trainable parameters. All model training and sampling were carried out using 8 NVIDIA H100 GPUs with Distributed Data Parallel. We trained the main ConfRover model for 180 epochs ($\sim 37$ hrs) and ConfRover–interp model for additional 220 epochs ($\sim 45$ hrs).
Inference cost varies with protein size and benchmark scale. For a moderately sized protein (e.g., 300 amino acids), simulating a short trajectory of 9 frames takes less than 1 minute, while generating an 80-frame trajectory (equivalent to 100 ns ATLAS setting in the paper) takes approximately 8.3 minutes. In comparison, a standard MD simulation under the ATLAS setup requires roughly 10 hours to reach 100 ns.
For our large-scale benchmarks, multi-start trajectory simulation (2,700+ trajectories) took 8 hours and 30 minutes, while time-independent sampling (250 conformations per protein) took 3 hours and 20 minutes.
A.3 Baseline Limitation
As an emerging research area, few models currently support learning protein dynamics in transferable settings. Existing approaches based on forward transport operators have been primarily trained and evaluated on small peptides (e.g., Timewarp [21]) or small fast-folding proteins (e.g., EquiJump [8]). Although AlphaFolding [6] was trained on ATLAS, its models are not yet publicly available. Due to these limitations, we use MDGen as the only available model for trajectory-based comparison. For conformation interpolation, neither sampling-based method [10] nor video-like method MDGen [18] have been trained and evaluated on large proteins.
Appendix B Diffusion Models on SE(3) Space
Diffusion Probabilistic Models (DPM) model complex distributions through iterative denoising. In the context of protein conformations, DPMs defined over $\mathrm{SE(3)}$ translation-rotation space have been applied for protein backbone structural generation [43, 40]. Following Section 2.2, $\mathbf{x}_{0}=(\mathbf{T}_{0},\mathbf{R}_{0})\in\mathrm{SE(3)}^{N}$ denotes the translations and rotations of backbone rigids in data. The diffusion processes defined in the translation and rotation subspace add noise to corrupt the data:
$\mathrm{d}\mathbf{T}_{t}$ $=-\frac{1}{2}\beta_{t}\mathrm{P}\mathbf{T}_{t}\mathrm{~{}d}t+\sqrt{\beta_{t}}\mathrm{P}\mathrm{d}\mathbf{w}_{t},$
$\mathrm{~{}d}\mathbf{R}_{t}$ $=\sqrt{\frac{\mathrm{~{}d}}{\mathrm{~{}d}t}\sigma^{2}_{t}}\mathrm{~{}d}\mathbf{w}_{t}^{\mathrm{SO(3)}},$
where $t\in[0,1]$ is the diffusion time, $\beta_{t}$ and $\sigma_{t}$ are predefined time-dependent noise schedules and $\mathrm{P}$ is a projection operator removing the center of mass. $\mathbf{w}_{t}$ and $\mathbf{w}_{t}^{\mathrm{SO(3)}}$ are the standard Wiener processes in $\mathcal{N}(0,I_{3})^{\otimes N}$ and $\mathcal{U}(\mathrm{SO}(3))^{\otimes N}$ respectively.
[…]
The associated reverse-time stochastic differential equation (SDE) follows:
$\mathrm{d}\mathbf{T}_{t}$ $=\mathbf{P}\left[-\frac{1}{2}\beta_{t}\mathbf{T}_{t}-\beta_{t}\nabla\log p_{t}(\mathbf{T}_{t})\right]\mathrm{d}t+\sqrt{\beta_{t}}\mathrm{P}\mathrm{d}\bar{\mathbf{w}}_{t},$
$\mathrm{d}\mathbf{R}_{t}$ $=-\frac{\mathrm{~{}d}}{\mathrm{~{}d}t}\sigma^{2}_{t}\nabla\log p_{t}(\mathbf{R}_{t})\mathrm{d}t+\sqrt{\frac{\mathrm{d}}{\mathrm{d}t}\sigma^{2}_{t}}\mathrm{d}\bar{\mathbf{w}}_{t}^{\mathrm{SO(3)}},$ (6)
where $\bar{\mathbf{w}}_{t}$ and $\bar{\mathbf{w}}_{t}^{\mathrm{SO(3)}}$ denote standard Wiener processes in the reverse time.
The reverse process can be approximated by a neural network through the denoising score matching loss for translation and rotation:
$\mathcal{L}(\theta)=$ $\mathcal{L}^{\mathbf{T}}(\theta)+\mathcal{L}^{\mathbf{R}}(\theta)$ (7)
$=$ $\mathbb{E}\left[\lambda(t)\|s_{\theta}(\mathbf{T}_{t},t)-\nabla_{\mathbf{T}_{t}}\log p_{t|0}(\mathbf{T}_{t}|\mathbf{T}_{0})\|^{2}\right]$
$+\mathbb{E}\left[\lambda^{r}(t)\|s_{\theta}^{r}(\mathbf{R}_{t},t)-\nabla_{\mathbf{R}_{t}}\log p_{t|0}(\mathbf{R}_{t}|\mathbf{R}_{0})\|^{2}\right],$
where $\lambda(t)$ and $\lambda^{r}(t)$ are time-dependent weights, $s_{\theta}(\mathbf{T_{t},t})$ and $s^{r}_{\theta}(\mathbf{R}_{t},t)$ are the score networks commonly parameterized with shared weights. The expectations are taken over diffusion time $t\sim\mathcal{U}[t_{\mathrm{min}},1]$, and over noisy and clean data pairs from the forward process $(\mathbf{T}_{0},\mathbf{T}_{t})$ and $(\mathbf{R}_{0},\mathbf{R}_{t})$.
Appendix C Derivation of Equations in Section 3.2
Here we provide a more rigorous derivation of equations in Section 3.2. For clarity, we omit the conditioning variable $\mathcal{P}$ in the intermediate steps.
As defined in Equation (1), our goal is to model the frame-level conditional distribution $p(\mathbf{x}^{l}|\mathbf{x}^{<l})$. To achieve this, we encode the previous frames $\{\mathbf{x}^{i}\}_{i=1}^{l-1}$ into a sequence of latent embeddings $\{\mathbf{h}^{i}\}_{i=1}^{l-1}$, and model inter-frame dependencies in this latent space.
By applying the Bayes’ rule, we can factor the joint distribution over the current conformation $\mathbf{x}^{l}$ and intermediate latent embeddings $\mathbf{h}=(\mathbf{h}^{l},\mathbf{h}^{<l})$ as:
$p(\mathbf{x}^{l},\mathbf{h}^{l},\mathbf{h}^{<l}\mid\mathbf{x}^{<l})=p(\mathbf{x}^{l}\mid\mathbf{h}^{l},\mathbf{h}^{<l},\mathbf{x}^{<l})\,p(\mathbf{h}^{l}\mid\mathbf{h}^{<l},\mathbf{x}^{<l})\,p(\mathbf{h}^{<l}\mid\mathbf{x}^{<l}).$
Integrating both sides over the latent variables $\mathbf{h}$ yields:
$p(\mathbf{x}^{l}\mid\mathbf{x}^{<l})=\int_{\mathbf{h}}p(\mathbf{x}^{l}\mid\mathbf{h}^{l},\mathbf{h}^{<l},\mathbf{x}^{<l})\,p(\mathbf{h}^{l}\mid\mathbf{h}^{<l},\mathbf{x}^{<l})\,p(\mathbf{h}^{<l}\mid\mathbf{x}^{<l})\,\mathrm{d}\mathbf{h}.$ (8)
In our approach, both $p(\mathbf{h}^{l}\mid\mathbf{h}^{<l},\mathbf{x}^{<l})$ and $p(\mathbf{h}^{<l}\mid\mathbf{x}^{<l})$ are modeled using deterministic neural networks: an encoder $f_{\eta}^{\text{enc}}(\mathbf{x}^{i})$ and an autoregressive temporal module $f_{\xi}^{\text{temp}}(\mathbf{h}^{<l})$, respectively.
These mappings reduce both $p(\mathbf{h}^{l}\mid\mathbf{h}^{<l},\mathbf{x}^{<l})$ and $p(\mathbf{h}^{<l}\mid\mathbf{x}^{<l})$ to Dirac delta functions, and the conditional dependencies can be simplified as:
$p(\mathbf{h}^{l}|\mathbf{h}^{<l},\mathbf{x}^{<l})=p(\mathbf{h}^{l}|\mathbf{h}^{<l})$
$p(\mathbf{x}^{l}|\mathbf{h}^{l},\mathbf{h}^{<l},\mathbf{x}^{<l})=p(\mathbf{x}^{l}|\mathbf{h}^{l}).$
Substituting into Equation (8) gives:
$p(\mathbf{x}^{l}|\mathbf{x}^{<l})$ $=\int_{\mathbf{h}}p(\mathbf{x}^{l}|\mathbf{h}^{l})\cdot p(\mathbf{h}^{l}|\mathbf{h}^{<l})\cdot p(\mathbf{h}^{<l}|\mathbf{x}^{<l})\mathrm{d}\mathbf{h}.$ (9)
Again, due to the deterministic nature of the encoder and temporal module, there is no marginalization involved in Equation (9), yielding:
$p(\mathbf{x}^{l}|\mathbf{x}^{<l})$ $=p(\mathbf{x}^{l}|\mathbf{h}^{l})$
$\text{where}\quad\mathbf{h}^{i}$ $=f_{\eta}^{\text{enc}}(\mathbf{x}^{i},\mathcal{P}),\quad i=1,2,\dots,l-1$
$\text{and}\quad\mathbf{h}^{l}$ $=f_{\xi}^{\text{temp}}\left(\mathbf{h}^{1},\mathbf{h}^{2},\dots,\mathbf{h}^{l-1}\right).$
Finally, we approximate $p(\mathbf{x}^{l}|\mathbf{h}^{l})$ with a parameterized model $p_{\theta}^{\text{enc}}(\mathbf{x}^{l}|\mathbf{h}^{l})$.
Appendix D Method Details
D.1 Detailed Module Architectures
Encoding Layer.
The protein-specific single and pair representations are obtained from the Evoformer stack of a pretrained OpenFold model (with frozen weights), after three recycle iterations. In addition, we encode residue-level sequence information by combining sinusoidal positional embeddings of residue indices with learnable embeddings for the 20 standard amino acid types. These features are concatenated with the single representation from the FoldingModule.
To encode the structural information of each conformation frame, we introduced the FrameEncoder, a pseudo-beta-carbon (Cβ) coordinate encoder similar to the InputEmbedding module from AlphaFlow [17] (without diffusion time embedding). Specifically, this module first compute the pairwise distances between residues using Cβ coordinates. These distances are then binned, embedded into latent embedding, and further refined through triangular update blocks including triangle attention and multiplication updates [19]. See Algorithm 1 for the specifics. The resulting per-frame Cβ pair embedding $\mathbf{z}_{\text{Frame}}^{l}$ is concatenated with the pair representation from FoldingModule.
Both single and pair embeddings are projected into the same dimension of $d$ for simplicity, forming the latent embedding $\mathbf{h}^{l}=[\mathbf{s}^{l},\mathbf{z}^{l}]$ for each frame. See detailed illustration in Figure 8.
Figure 8: Architecture details of the Encoding Layer. A frozen FoldingModule encodes the protein-specific information from its sequence, containing prior knowledge on its chemical environment and folding structures. The single representation is further concatenated with additional amino acid embeddings and projected to a hidden dimension of size $d$; The pair representation is concatenated with frame conformation information, encoded in Cβ pair embedding, and projected to a hidden dimension of size $d$. Both frame-level single and pair embeddings form the frame-level latent for downstream modules.
Algorithm 1 FrameEncoder
  Input: Pseudo beta carbon (Cβ) coordinates $\mathbf{x}\in\mathbb{R}^{N\times 3}$, time $t\in[0,1]$
  Output: Input pair embedding $\mathbf{z}\in\mathbb{R}^{N\times N\times 64}$
  $\mathbf{z}_{ij}\leftarrow\lVert\mathbf{x}_{i}-\mathbf{x}_{j}\rVert$
  $\mathbf{z}_{ij}\leftarrow\operatorname{Bin}(\mathbf{z}_{ij},{\min}=3.25\text{ \AA},{\max}=50.75\text{ \AA},N_{\text{bins}}=39)$
  $\mathbf{z}_{ij}\leftarrow\operatorname{Linear}(\operatorname{OneHot}(\mathbf{z}_{ij}))$
  for $l\leftarrow 1$ to $N_{\text{blocks}}=4$ do
     $\{\mathbf{z}\}_{ij}\mathrel{+}=\operatorname{TriangleAttentionStartingNode}({\mathbf{z}_{ij}},c=64,N_{\text{head}}=4)$
     $\{\mathbf{z}\}_{ij}\mathrel{+}=\operatorname{TriangleAttentionEndingNode}({\mathbf{z}_{ij}},c=64,N_{\text{head}}=4))$
     $\{\mathbf{z}\}_{ij}\mathrel{+}=\operatorname{TriangleMultiplicationOutgoing}({\mathbf{z}_{ij}},c=64)$
     $\{\mathbf{z}\}_{ij}\mathrel{+}=\operatorname{TriangleMultiplicationIncoming}({\mathbf{z}_{ij}},c=64)$
     $\{\mathbf{z}\}_{ij}\mathrel{+}=\operatorname{PairTransition}({\mathbf{z}_{ij}},n=2)$
  end for
  $\mathbf{z}_{ij}=\operatorname{LayerNorm}(\mathbf{z}_{ij})$
Trajectory Module.
In the Trajectory Module, we interleave layers of StructuralUpdate and TemporalUpdate to iteratively update the latent $[\mathbf{s}^{l},\mathbf{z}^{l}]$, enabling the temporal reasoning across frames and structural refinement within each frame.
For the StructuralUpdate, we adopt a Pairformer block from AlphaFold 3 [2], which jointly updates the single and pair embeddings of the current frame through structural reasoning. After the StructuralUpdate, the pair embedding is flattened from $[N,N,d]$ to $[N\times N,d]$ and concatenated with the single embedding before being passed into the TemporalUpdate.
To model temporal dependencies between frames, we use a lightweight Llama architecture [38]. We transpose the input such that the temporal dimension is treated as the sequence axis for channel-wise self-attention across time. Rotary positional encoding [37] is applied to encode the temporal position for each frame. A causal attention mask is applied to restrict each frame to only attend to previous frames. After the temporal update, the latent embedding are reshaped and split back into single and pair embeddings.
Figure 9 and Table 5 provide the detailed module architecture and hyperparameter configurations, respectively. A StructuralUpdate block is included for every two TemporalUpdate layers.
Figure 9: Architecture details of the Trajectory Module. Trajectory Module contains interleaving blocks of StructuralUpdate and TemporalUpdate (only one block of each is shown). StructuralUpdate leverages the Pairformer architecture from Abramson et al. [1], updating the pair embeddings with triangular updates and the single embeddings using with pair bias from the updated pair. The updated pair embeddings are flatten and concatenate with single embedding for channel-wise temporal update. The attention is applied along the temporal dimension and update each single and pair embedding channels independently. The embeddings from TemporalUpdate are split and reshape back into single and pair embeddings.
Table 5: Hyperparameter choices of Trajectory Module
              Hyperparameters               Values
              Lightweight Llama (TemporalUpdate)
              Number of layers               8
              Dimension of the MLP embeddings               256
              Dimension of the hidden embeddings               128
              Number of attention heads               4
              Pairformer (StructuralUpdate)
              Dimension of single embeddings               128
              Dimension of pair embeddings               128
              Dimension of triangle multiplication hidden embeddings               128
              Number of triangle attention heads               4
              Dimension of pair attention embeddings               32
              Transition layer expanding factor               4
              Pair attention dropout rate               0.25
Structure Decoder.
Conformation generation, conditioned on the temporal signals from the Trajectory Module, is performed using the model architecture from ConfDiff [40], As shown in Figure 10, following ConfDiff, the inputs to the denoising model include diffusion time $t$, pairwise distance between residue rigids, and residue indices (not shown). They are encoded and concatenated with the single and pair embeddings from FoldingModule. In addition, the single and pair embeddings from the Trajectory Module are projected back to the latent dimension $d_{\mathbf{s}}$ and $d_{\mathbf{z}}$, respectively, to modify the single and pair embeddings used by ConfDiff.
The core of the Structure Decoder consists of multiple IPA-transformer blocks, which update single and pair embeddings as well as the SE(3) rigids of noisy conformations. In the final block, torsional angles are predicted by TorsionPred and, together with denoised rigids, to reconstruct the atomic structure of generated conformation. The corresponding hyperparameter settings are summarized in Table 6.
Figure 10: Architecture details of the Structure Decoder. Single and pair embeddings from Trajectory Module is used to update the original embeddings from FoldingModule. The resulting embeddings are fed into blocks of IPA-transformer to update single, pair embeddings and denoise rigids, SE(3) representation of protein backbone conformations. Denoised rigids together with torsion angles predicted by TorsionPred recovers the atomic structure of protein conformation at this frame.
Table 6: Hyperparameter choices of the Structure Decoder
              Hyperparameters               Values
              Neural network
              Number of IPA blocks               4
              Dimension of single embedding ($d_{\mathbf{s}}$)               256
              Dimension of pair embedding ($d_{\mathbf{z}}$)               128
              Dimension IPA hidden embedding               256
              Number of IPA attention heads               4
              Number of IPA query points               8
              Number of IPA value points               12
              Number of transformer attention heads               4
              Number of transformer layers               2
              SE(3)-diffusion SDE
              Number of diffusion steps               200
              Translation scheduler               Linear
              Translation $\beta_{\min}$               0.1
              Translation $\beta_{\max}$               20
              Rotation scheduler               Logarithmic
              Rotation $\sigma_{\min}$               0.1
              Rotation $\sigma_{\max}$               1.5
D.2 Training and Inference details
We train all ConfRover models on the trajectories from the ATLAS training set, following the train-validation-test split of previous works [17, 40, 18]. Specifically, we exclude the training proteins longer than 384 amino acid residues, leading to 1080 training proteins.
Most components of ConfRover models were trained from scratch, except for the FoldingModule, where we used frozen weights from OpenFold to extract the single and pair representations from three recycling iterations, and the DiffusionDecoder, which was initialized from ConfDiff-OF-r3-MD checkpoint provided by the authors [^1] https://github.com/bytedance/ConfDiff.
During each training epoch, we randomly sample stride length from $2^{0}$ to $2^{10}$ to extract sub-trajectories of length $L=9$ at varying time scales. With the use of causal transformers, input frames were shifted forward by one frame with a [MASK] token padded at the beginning of the trajectory. Combined with the use of a causal mask in temporal attention, the design ensures that each frame is trained to sample conditioned only on previous frames and the first frame is generated unconditionally using only the [MASK] token as input.
We trained main ConfRover model 180 epochs ($\sim 37$ hrs) and ConfRover–interp model for additional 220 epochs ($\sim 45$ hrs). Additional training hyperparameters can be found in Table 7. All model training and sampling were carried out using 8 NVIDIA H100 GPUs.
Table 7: Training hyperparameters
              Hyperparameters               Values
              Batch Size               1
              Frames Num               8
              Gradient Clip               1.0
              Learning Rate               1×$\times 10^{-4}$
              Optimizer               Adam (weight decay = 0.)
Appendix E Additional Experimental Results
E.1 Trajectory Simulation: multi-start
Benchmark Curation.
In multi-start, we sample short trajectories from varying starting point while ensuring the generation within the scope of the reference trajectory. For example, we select frame index of 1000, 3000, 5000, and 7000 as starting frames for stride $S=128/256$, resulting in 12 test trajectories from triplicates; frame index 1000, 3000, 5000 as starting frames for stride $S=512$, and frame index 1000 for stride $S=1024$ to avoid exceeding total of 10000 frames. This provided us 2,706 different starting conditions from 82 proteins from the ATLAS test set for evaluation.
PCA Projection.
[…]
Additional Results.
Here we also include the scatterplot of Pearson correlation in Figure 11 and additional metrics from the multi-start experiments at different strides in Table 8. Across different strides, ConfRover models consistently outperforms MDGen in recovering the correct level of dynamics.
Figure 11: Scatterplots of conformation changes in sample trajectories versus those in the reference trajectories, measured by trajectory-level conformation changes, frame-level conformation changes, and next-frame difference ($\Delta$Frame), measured by the RMSD of alpha carbons (unit: Å) or $L^{2}$ distance in the projected PCA space. MDGen tends underestimate the magnitude of conformation changes while ConfRover generate samples at similar level as the MD reference. The exact match of measured conformation changes is not possible due to stochastic sampling in both MD simulation and generative models.
Table 8: Additional metrics from the multi-start benchmark. Results for different strides are shown in separate blocks. The better score in each block is highlighted in bold (excluding diversity)
Diversity MAE on PCA-2D ($\downarrow$) MAE on CA coordinates ($\downarrow$) Quality
Pairwise
RMSD
Trajectory Frame
Frame
Next
Trajectory Frame
Frame
Next
PepBond
Break %
($\downarrow$)
Stride=128
MDGen 1.26 6.53 1.28 0.91 5.02 0.96 0.71 27.9
ConfRover 1.63 3.10 1.10 0.74 3.83 0.81 0.64 17.3
Stride=256
MDGen 1.34 7.66 1.54 1.06 6.00 1.14 0.85 27.9
ConfRover 1.78 3.91 1.28 0.87 4.60 0.91 0.75 16.6
Stride=512
MDGen 1.40 9.12 1.94 1.25 7.27 1.41 1.01 28.0
ConfRover 1.89 4.84 1.53 1.01 5.66 1.07 0.89 16.7
Stride=1024
MDGen 1.51 11.48 2.62 1.55 9.04 1.80 1.24 28.1
ConfRover 2.04 6.75 1.89 1.25 7.39 1.26 1.09 16.7
Additional Visualization of Trajectory.
We additional unfiltered examples (randomly selected) for visual comparison of conformations generated by different models in the multi-start experiments, as shown in Figure 12.
Figure 12: Visualization of 10 trajectories randomly selected from the Multi-start benchmark. Trajectory conformations are colored by their secondary structures and superposed to show the dynamic ensemble. MDGen exhibits primarily local motions while ConfRover better reflects the motions observed in MD reference.
E.2 Trajectory Simulation: 100 ns Long Trajectory Simulation on ATLAS
Details on Evaluation Metrics.
[…]
Additional Visualizations.
We additionally provide unfiltered (randomly selected) PCA plots in Figure 13.
Figure 13: Visualization the ATLAS-100ns trajectories from 20 randomly selected cases. The blue background indicates the density of the ground-truth conformation distribution from MD reference. ConfRover shows improved conformation state recovery in several cases (e.g., 7JRQ-A, 6YHU-B, 7AQX-A, 6L4L-A, 6OZ1-A, etc), sampling more diverse conformations. Yet, the gap between the oracle MD 100ns and deep learning models is evident in some cases (e.g., 7JFL-C, 6TLY-A, 6JWH-A, etc)
E.3 Time-independent Conformation Sampling
We follow the evaluation protocol of Ye et al. [42] to assess time-independent conformation sampling on the ATLAS test set. For each protein, 250 independent samples are generated. Since MDGen does not support time-independent sampling, we approximate its performance using samples from its 100-ns trajectory, serving as a sequential-sampling baseline. The performance of state-of-the-art models, AlphaFlow and ConfDiff, is taken from Table 10 of Ye et al. [42], using their best-performing variants: AlphaFlow-MD and ConfDiff-Open-MD. Full results are shown in Table 9.
Table 9: Performance on time-independent conformation generation on ATLAS. A total of 250 conformations were sampled for each protein, and the median values across 82 proteins are reported. The best performance is highlighted in bold, and the second-best is underlined.*MDGen does not support time-independent sampling and the result from sequential sampling is used.
Diversity Flexibility: Pearson $r$ on Distributional accuracy
Pairwise
RMSD
RMSF
Pairwise
RMSD ↑
Global
RMSF ↑
Per target
RMSF ↑
RMWD ↓
MD PCA
$\mathcal{W}_{2}$
Joint
PCA $\mathcal{W}_{2}$
PC sim
$>$ 0.5 %↑
AlphaFlow 2.87 1.63 0.53 0.66 0.85 2.64 1.55 2.29 39.0
ConfDiff 3.43 2.21 0.59 0.67 0.85 2.75 1.41 2.27 35.4
MDGen* 1.38 0.78 0.41 0.49 0.74 2.81 1.95 2.38 13.4
ConfRover-traj 3.21 1.74 0.47 0.61 0.85 2.83 1.41 2.30 41.5
ConfRover 3.63 2.23 0.51 0.64 0.85 2.62 1.39 2.28 36.6
Ensemble observables Quality
Weak
contacts $J$
Transient
contacts $J$
Exposed
residue $J$
Exposed MI
matrix $\rho$
CA clash
PepBond
break % ↓
AlphaFlow 0.62 0.41 0.69 0.35 0.0 22.2
ConfDiff 0.63 0.39 0.65 0.33 0.5 6.5
MDGen* 0.50 0.28 0.57 0.26 0.3 29.0
ConfRover-traj 0.53 0.36 0.60 0.27 0.4 12.2
ConfRover 0.61 0.38 0.67 0.32 0.5 19.2
E.4 Conformation Interpolation
In conformation interpolation experiment, we selected trajectories from multi-start. These trajectories exhibit sufficient conformation changes (e.g., RMSD between the start and end frames $>4$) and clear interpolation path in the PCA space. The list of selected cases are in Table 10.
[…]
0 5ZNJ-A-R2F1000S512 6L4L-A-R1F3000S256 6TGK-C-R1F5000S256 7JFL-C-R2F3000S256
1 6E7E-A-R3F3000S512 6LRD-A-R1F1000S1024 6TLY-A-R2F1000S256 7JRQ-A-R1F1000S1024
2 6GUS-A-R2F7000S256 6LUS-A-R2F3000S128 6XDS-A-R1F1000S128 7LA6-A-R1F1000S512
3 6H49-A-R2F1000S1024 6OVK-R-R2F7000S128 6XRX-A-R3F5000S128 7LP1-A-R2F5000S256
4 6H86-A-R2F1000S1024 6OZ1-A-R1F1000S512 6Y2X-A-R2F3000S512 7P41-D-R2F3000S256
5 6IN7-A-R3F5000S128 6P5H-A-R1F1000S1024 7AEX-A-R3F5000S256 7P46-A-R3F5000S256
6 6J56-A-R1F3000S256 6P5X-B-R1F3000S256 7AQX-A-R2F3000S512 7RM7-A-R3F3000S128
7 6JPT-A-R2F5000S512 6Q9C-A-R3F3000S256 7ASG-A-R1F7000S128 7S86-A-R3F5000S256
8 6JV8-A-R3F1000S256 6QJ0-A-R1F1000S256 7BWF-B-R3F3000S128
9 6KTY-A-R3F1000S1024 6RRV-A-R1F1000S256 7C45-A-R1F5000S512
Figure 14: Normalized PCA distance of intermediate frames to the start and end frames, averaged over 38 cases selected from the multi-start benchmark. [Left] a comparison between ConfRover-interp and ConfRover, where ConfRover-interp generates smooth pathways while ConfRover does not; [Right] a comparison between ConfRover-interp and reference trajectories.
Additional Visualizations.
Here we include additional visualization on interpolation results.
Figure 15: Example interpolations results. ConfRover-interp generates smooth pathways between the start and end frames, capturing the dynamics observed with the MD reference while ConfRover does not show the correct intermediate conformations. Start and end frames are shown as solid structures; intermediate conformations are shown in fading colors. Main motions are indicated by blue dashed arrows. These examples highlight the difference between the original ConfRover and ConfRover-interp that further trained on the interpolation objective. The original ConfRover can miss key motions of the transition while ConfRover-interp correctly capture these motions.
Figure 16: PCA plot of 38 selected interpolation cases. MD reference trajectories and results from ConfRover-interp are shown in each plot.
E.5 Retraining MDGen for Experiment-Specific Setups
Due to its non-autoregressive design and training on fixed-length trajectories, MDGen cannot directly generate sequences of varying lengths. Therefore, in our evaluation, we generate trajectories using the original settings (named MDGen-SstrideFlength) and subsample them to match the evaluation setup. However, this post-processing may introduce artifacts. To address this concern, we retrain MDGen under the evaluation settings and compare the results on the Multi-start benchmark (stride = 256) and 100 ns long trajectory simulations, as shown below in Table 11, Table 12, Table 13 and Figure 17. As shown in these experiments, no significant difference of performance for key metrics observed comparing MDGen with post-processed results and models specifically trained at the evaluation settings, suggesting no evident decrease of trajectory quality from the subsampling post-process.
Table 11: Compare MDGen-S256F9 with MDGen from subsampling post-process. Here is the table summarizing the Pearson correlations of conformation changes between sampled and reference trajectories in multi-start. MDGen-S256F9 is trained and sampled with stride of 256 MD snapshots and length of 9 frames. The best scores are highlighted in bold.
[…]
Trajectory Frame $\Delta$Frame
MDGen 0.57 0.46 0.41
MDGen-S256F9 0.56 0.45 0.38
ConfRover 0.77 0.62 0.53
PCA 2D
Trajectory Frame $\Delta$Frame
MDGen 0.18 0.13 0.11
MDGen-S256F9 0.21 0.19 0.11
ConfRover 0.75 0.5 0.44
Table 12: Compare MDGen-S256F9 with MDGen from subsampling post-process. Here is the table summarizing additional metrics in multi-start benchmark. MDGen-S256F9 is trained and sampled with stride of 256 MD snapshots and length of 9 frames. The best scores are highlighted in bold.
Diversity MAE on PCA-2D ($\downarrow$) MAE on CA coordinates ($\downarrow$) Quality
Pairwise
RMSD
Traj. Frame $\Delta$Frame Traj. Frame $\Delta$Frame
PepBond
Break %($\downarrow$)
MDGen 1.34 7.66 1.54 1.06 6.00 1.14 0.85 27.9
MDGen-S256F9 1.59 6.90 1.47 1.03 5.33 1.13 0.82 16.2
ConfRover 1.78 3.91 1.28 0.87 4.60 0.91 0.75 16.6
Table 13: Compare MDGen-S256F9 with MDGen from subsampling post-process. Here is the table summarizing the state recovery performance in 100 ns long trajectory simulation. MDGen-S120F80 is trained and sampled with stride of 120 MD snapshots and length of 80 frames. The best scores are highlighted in bold. MD 100ns is included as the oracle.
JSD ($\downarrow$) Recall ($\uparrow$) F1 ($\uparrow$)
MD 100ns 0.31 0.67 0.79
MDGen 0.56 0.30 0.44
MDGen-S120F80 0.57 0.29 0.42
ConfRover 0.51 0.42 0.58
Figure 17: Compare MDGen-S256F9 with MDGen from subsampling post-process. This figure shows correlations of main dynamic modes between sampled trajectory and reference trajectory. MDGen-S120F80 is trained and sampled with stride of 120 MD snapshots and length of 80 frames.

### v2

@@
 ConfRover is the first model to sample both protein conformations and trajectories within a single framework, offering a novel and flexible approach for learning from protein MD data.
-00footnotetext: Equal contribution. Work done during their internship at ByteDance. Correspondence to Quanquan Gu <quanquan.gu@bytedance.com>.
+Project website: https://bytedance-seed.github.io/ConfRover.
+00footnotetext: Equal contribution. Work done during their internship at ByteDance. Corresponding Author.
 1 Introduction
@@
 Their ability to transition between different conformations enables biological processes critical to life.
-Characterizing the behavior of a protein, including its (1) dynamic motions, (2) conformational distribution, and (3) transitions between different states, is crucial for understanding its function and guiding the design of novel proteins [4, 11, 31].
-Molecular dynamics (MD) simulations are the “gold standard” for studying protein conformational changes [7, 20, 32].
+Characterizing the behavior of a protein, including its (1) dynamic motions, (2) conformational distribution, and (3) transitions between different states, is crucial for understanding its function and guiding the design of novel proteins [berendsen2000collectivedynamics, frauenfelder1991energydynamics, mccammon1984proteindynamics].
+Molecular dynamics (MD) simulations are the “gold standard” for studying protein conformational changes [Childers2017_mddesign, karplus2005molecularmd, mccammon1977dynamicsmdbpti].
 These simulations use physical models to describe the energy of a protein conformation and the forces acting on its atoms.
-By simulating atomic motion through classical mechanics and iteratively sampling conformations over time, MD enables researchers to explore different conformations, approximating the conformational distribution at equilibrium, and gain mechanistic insights into protein’s behavior.
+By simulating atomic motion through classical mechanics and iteratively sampling conformations over time, MD enables researchers to explore different conformations, approximating the conformational distribution at equilibrium, and gain mechanistic insights in protein functions.
 However, MD simulations are both computationally expensive and technically challenging due to long simulation times and the tendency to become trapped in local energy minima.
 Figure 1: Key ideas of ConfRover.
-(A) Various conformation generation tasks are unified as conditional generation, learned via frame-level autoregressive factorization.
-Each block represents a frame and conditioning frames are outlined in black.
-The arrow indicates frame dependency.
+(A) Conformation generation tasks with various conditioning configurations.
+Each block denotes a frame and arrows indicates the sequential dependencies among frames from autoregressive formulation.
+Initial conditioning frames are outlined in black.
+In conformation interpolation, the last frame is repositioned and prepended to the first frame for proper sequential dependencies.
 (B) ConfRover models each frame as a conditional distribution given preceding frames.
 Sequential dependencies are captured through latent variables $\mathbf{h}$, and conformations are sampled from a diffusion decoder, conditioned on the updated latent.
-These challenges have motivated the use of deep generative models to study proteins, leveraging the rich conformational and dynamic information provided by large-scale MD datasets [26, 39]:
+These challenges have motivated the use of deep generative models to study proteins, leveraging the rich conformational and dynamic information provided by large-scale MD datasets [liuDynamicPDBNew2024, yannATLASProteinFlexibility2024]:
 (1) Generating the dynamic motions of proteins is a direct analog to MD simulation.
-Pioneering works modeled this by learning transition probabilities of future conformations from the current state [21, 34, 8].
+Pioneering works modeled this by learning transition probabilities of future conformations from the current state [costa2024equijump, klein2024timewarp, schreiner2024implicit].
 However, MD trajectories are often non-Markovian due to partially observed coordinates (e.g., protein-only atoms) and environmental coupling (e.g., using Langevin thermostats).
-To mitigate this, Cheng et al.
-[6] incorporated higher-order information using multiple context frames, though this requires fix context windows and limits flexibility.
-Jing et al.
-[18] instead modeled the joint distribution over the entire trajectory, capturing complex dependencies among frames.
+To mitigate this, cheng20244d4ddiff incorporated higher-order information using multiple context frames, though this requires fix context windows and limits flexibility.
+jing2024generative_mdgen instead modeled the joint distribution over the entire trajectory, capturing complex dependencies among frames.
 Due to training on fixed-length trajectories and the non-autoregressive design, their model has limited inference-time flexibility that cannot generate variable-length trajectories.
-Li et al.
-[25] introduced an autoregressive approach for flexible trajectory extension, but its deterministic formulation cannot capture trajectory distributions or generate diverse samples.
+li2025geometricGST introduced an autoregressive approach for flexible trajectory extension, but its deterministic formulation cannot capture trajectory distributions or generate diverse samples.
 (2) Learning the conformational distribution enables sampling time-independent conformations.
-Several methods [33, 17, 40, 44, 23] train diffusion- or flow-based generative models on conformation ensembles from MD simulation data, bypassing the need for sequential sampling.
+Several methods [jingAlphaFoldMeetsFlow2024, lewis2024scalableBioemu, noe2019boltzmann, wang2024proteinconfdiff, zhengPredictingEquilibriumDistributions2024aDiG] train diffusion- or flow-based generative models on conformation ensembles from MD simulation data, bypassing the need for sequential sampling.
 While effective for generating samples in parallel, they disregard temporal information in MD trajectories and therefore cannot simulate physical motion of proteins.
 (3) Conformation interpolation generate transition pathways between different states.
-Recent works [10, 18] have extended generative modeling to conformation interpolation, where the goal is to generate plausible intermediate samples between known start and end conformational states.
-Jing et al.
-[18] framed interpolation as a conditional trajectory generation task, but it requires training task-specific model and has not been evaluated on large proteins.
+Recent works [du2024doobs, jing2024generative_mdgen] have extended generative modeling to conformation interpolation, where the goal is to generate plausible intermediate samples between known start and end conformational states.
+jing2024generative_mdgen framed interpolation as a conditional trajectory generation task, but it requires training task-specific model and has not been evaluated on large proteins.
 As these generative problems all stem from the same underlying physical principles and involve sampling from a protein’s conformational space, a natural question arises: can we develop a general framework to learn all of these objectives?
@@
   • We introduce a simple yet general framework to learn both the conformational distribution and dynamics from MD data, supporting multiple generation tasks including trajectory simulation, time-independent conformation sampling, and conformation interpolation.
-  • We design a modular architecture that captures temporal dependencies in latent space using efficient causal transformers (i.e., Llama [38]), and directly models conformations in continuous SE(3) space using a diffusion decoder, avoiding discretization structure into tokens.
-  • Experiments show strong capabilities of ConfRover: it outperforms MDGen [18] in trajectory simulation, matches the performance of AlphaFlow [17] and ConfDiff [40] in time-dependent generation, and can effectively sample conformations interpolating two endpoints.
+  • We design a modular architecture that captures temporal dependencies in latent space using efficient causal transformers (i.e., Llama [touvron2023llama]), and directly models conformations in continuous SE(3) space using a diffusion decoder, avoiding discretization structure into tokens.
+  • Experiments show strong capabilities of ConfRover: it outperforms MDGen [jing2024generative_mdgen] in trajectory simulation, matches the performance of AlphaFlow[jingAlphaFoldMeetsFlow2024] and ConfDiff[wang2024proteinconfdiff] in time-dependent generation, and can effectively sample conformations interpolating two endpoints.
 2 Background
 2.1 Data Generation from Molecular Dynamics
-Molecular dynamics describes the motion of molecules through Newtonian mechanics $M\ddot{\mathbf{x}}=-\nabla U(\mathbf{x})$, where $\mathbf{x}$ denotes coordinates of atoms in the system, $M$ is the atomic mass, $U(\mathbf{x})$ is the potential energy of the configuration and $-\nabla U(\mathbf{x})$ represents the forces acting on atoms.
+Molecular dynamics describes the motion of molecules through Newtonian mechanics $M\ddot{\mathbf{x}}=-\nabla U(\mathbf{x})$, where $\mathbf{x}$ denotes coordinates of atoms in the system, $M$ is the atomic mass, $U(\mathbf{x})$ is the potential energy of the configuration and $\nabla U(\mathbf{x})$ represents the forces acting on atoms.
 In practice, stochastic and frictional forces are integrated to model energy exchange with the environment and maintain temperature control of the system, converting the equations of motion to a Langevin process:
@@
 Over time, the samples converges to the Boltzmann distribution $p(\mathbf{x})\propto\exp(-U(\mathbf{x})/k_{B}T)$.
-The trajectory of protein coordinates, $\mathbf{x}^{1:L}_{\text{prot}}=(\mathbf{x}_{\text{prot}}^{1},\mathbf{x}_{\text{prot}}^{2},\dots,\mathbf{x}_{\text{prot}}^{L})$, is extracted and saved at prescribed simulation intervals, providing both distributional and kinetic information the protein conformational dynamics.
+The trajectory of protein coordinates, $\mathbf{x}^{1:L}_{\text{prot}}=(\mathbf{x}_{\text{prot}}^{1},\mathbf{x}_{\text{prot}}^{2},\dots,\mathbf{x}_{\text{prot}}^{L})$, is extracted and saved at prescribed simulation intervals, providing both distributional and kinetic information in the protein conformational dynamics.
 For simplicity, we omit the subscript ‘prot’ and use $\mathbf{x}$ to denote protein coordinates throughout the paper.
 2.2 Protein Representations
-[…]
+Proteins are chain-like molecules composed of amino acid residues, each selected from 20 standard amino acid types.
+We parameterize the coordinates of heavy atoms in a protein using the SE(3)-torsional convention [jumperHighlyAccurateProtein2021AF2]: the backbone atoms (N-$\alpha-C$) of each residue define a local coordinate via a Gram-Schmidt process, referred to as a rigid.
+The position and orientation of each rigid relative to the global coordinate system are described by a translation-rotation transformation in SE(3) space.
+The backbone conformation of a protein with $N$ residues can then be represented as $\mathbf{x}=(\mathbf{T},\mathbf{R})\in\mathrm{SE(3)}^{N}$, where $\mathbf{T}\in\mathbb{R}^{N\times 3}$ and $\mathbf{R}\in\mathrm{SO(3)}^{N}$ are the translation and rotation components.
+The coordinates of the oxygen atom of the backbone and the side chain atoms can be determined with the addition of up to 7 torsional angles $(\phi,\psi,\omega,\chi_{1},\dots,\chi_{4})$ describing the bond rotation.
+Therefore, the complete configuration of a protein structure is parameterized in the space: $\mathbf{x}=(\mathbf{T},\mathbf{R},\phi,\psi,\omega,\chi_{1},\dots,\chi_{4})\in(\mathrm{SE(3)}\times\mathbb{T}^{7})^{N}.$
 2.3 SE(3)-Diffusion Models for Protein Conformation Generation
 Diffusion generative models are capable of learning complex data distributions.
-Training involves progressively corrupting data with noise and learning to reverse this process through denoising, thereby modeling the original data distribution [13, 35].
-Recently, diffusion models operating in SE(3) space have been proposed to model protein backbone structures [43, 40].
+Training involves progressively corrupting data with noise and learning to reverse this process through denoising, thereby modeling the original data distribution [hoDenoisingDiffusionProbabilistic2020ddpm, songScoreBasedGenerativeModeling2021].
+Recently, diffusion models operating in SE(3) space have been proposed to model protein backbone structures [wang2024proteinconfdiff, yimSEDiffusionModel2023].
 Below, we provide a brief overview of diffusion model and defer the detailed SE(3) formulation to Appendix B:
@@
 Despite its simplicity, this formulation naturally supports multiple learning objectives in protein conformation modeling.
-In its base form, it models temporal dependencies across the frames, learning to generate the trajectory.
+In its base form, it models temporal dependencies among frames by learning to generate the trajectory.
 When $L=1$, it removes the frame context and reduces to a single-frame distribution $p(\mathbf{x}|\mathcal{P})$, learning to direct sample time-independent conformations.
@@
 By prepending conditioning frames $\mathcal{K}$ to the sequence, we train the model to learn any conditional generation $p(\mathbf{x}^{1:L}|\mathcal{K},\mathcal{P})$, including conformation interpolation by setting $\mathcal{K}=\{\mathbf{x}^{1},\mathbf{x}^{L}\}$.
-Similar idea was applied in text infilling tasks by shuffling the order of text contexts [3].
+A similar idea was applied in text infilling tasks by shuffling the order of text contexts [bavarian2022efficienttraininglanguagemodelsInfill].
 After defining the main learning objectives for trajectory simulation, single-frame (time-independent), and conformation interpolation, we describe how to effectively model autoregressive dependencies over protein conformations in Section 3.2.
@@
 We propose a modular design composed of an encoder, a latent sequence model, and a stochastic decoder.
-This design enables the use of modern causal transformers, such as in Llama [38], to efficiently capture sequential dependencies between frames in the latent space (Figure 2).
+This design enables the use of modern causal transformers, such as in Llama [touvron2023llama], to efficiently capture sequential dependencies between frames in the latent space (Figure 2).
 During training, the input sequence is shifted by one frame and a mask token “[M]” is prepended; the generation process also begins with the mask token and conditioning frames (e.g., $\mathbf{x}^{1}$).
@@
 $\mathbf{h}^{l}_{\text{updated}}$ $=f_{\xi}^{\text{temp}}\left(\mathbf{h}^{1},\mathbf{h}^{2},\dots,\mathbf{h}^{l-1}\right).$ (3)
-Since that both the encoder and the temporal module are chosen to be deterministic, $p(\mathbf{x}^{l}|\mathbf{x}^{<l})$ reduces to a conditional generation over the updated latent, realized through a probabilistic decoder:
+Since both the encoder and the temporal module are chosen to be deterministic, $p(\mathbf{x}^{l}|\mathbf{x}^{<l})$ reduces to a conditional generation over the updated latent, realized through a probabilistic decoder:
 $p(\mathbf{x}^{l}|\mathbf{x}^{<l},\mathcal{P})$ $=p_{\theta}^{\text{dec}}(\mathbf{x}^{l}|\mathbf{h}^{l}_{\text{updated}}).$ (4)
-Details on this latent modeling is provided in Appendix C.
+Details on this latent modeling are provided in Appendix C.
 As a result, we train a model $(\eta,\xi,\theta)$ by jointly optimizing the three modules $f_{\eta}^{\text{enc}}$, $f_{\xi}^{\text{temp}}$, and $p_{\theta}^{\text{dec}}$.
@@
 This trains the model to directly sample conformations.
-A similar strategy has been used in image-video training [27, 15].
+A similar strategy has been used in image-video training [ho_video_2022, liu2024mardini].
 3.3 Training Autoregressive Model with SE(3) Diffusion Loss
 The main challenge in applying autoregressive modeling to conformation trajectories lies in representing the continuous distribution of protein conformations within a framework typically used for discrete token sequences.
-While some studies have approached this by discretizing the protein structural space into discrete tokens [12, 28, 29], such approaches inherently suffer from discretization error, which can lead to suboptimal performance in modeling protein conformations.
-Instead, we propose to directly model the continuous conformational space using diffusion probabilistic models and employ the DSM loss for autoregressive model training, similar to Li et al.
-[24].
+While some studies have approached this by discretizing the protein structural space into discrete tokens [hayesSimulating500Million, liu_diffusion_2023pvqd, lu2024structureesmdiff], such approaches inherently suffer from discretization error, which can lead to suboptimal performance in modeling protein conformations.
+Instead, we propose to directly model the continuous conformational space using diffusion probabilistic models and employ the DSM loss for autoregressive model training, similar to li2024autoregressivemar.
 Specifically, we perform DSM loss training in SE(3) space.
@@
 Figure 3: Architecture overview.
-(A) Encoding Layer embeds protein sequence and input structure to each frame as a frame latent representation $\mathbf{h}^{l}$, comprised of single and pair embeddings; (B) The Trajectory Module then updates frame latent $\mathbf{h}^{l}$ using interleaved structural and temporal update blocks; (C) A diffusion-based Structure Decoder learns to denoise a noisy conformation conditioned on the updated frame latent $\mathbf{h}^{l}$; during inference, it samples conformations from the prior distribution.
+(A) Encoding Layer embeds protein sequence and input structure to each frame as a frame latent representation $\mathbf{h}^{l}$, comprised of single and pair embeddings; (B) The Trajectory Module then updates frame latent $\mathbf{h}^{l}$ using interleaved structural and temporal update blocks; (C) A diffusion-based Structure Decoder learns to denoise noisy conformations conditioned on the updated frame latent $\mathbf{h}^{l}$; during inference, it samples conformations from the prior distribution.
 See Appendix D.1 for details.
 Encoding Layer.
-A FoldingModule, parameterized by a pretrained OpenFold model [2], extracts protein-specific embeddings $\mathcal{P}$ consisting of a single representation ($\mathbf{s}$) and a pair representation ($\mathbf{z}$), shared across frames.
-For each frame, a FrameEncoder, adapted from the template module used in prior works [19, 17], encodes pairwise distance of pseudo–Cβ atoms via triangular updates and merges this frame pair representations $\mathbf{z}_{\text{Frame}}^{l}$ with the protein pair representation $\mathbf{z}$.
-The resulting frame latent embedding, $\mathbf{h}^{l}=[\mathbf{s}^{l},\mathbf{z}^{l}]$, are invariant to global translation and rotation of the conformations, and are passed to the Trajectory Module.
+A FoldingModule, parameterized by a pretrained OpenFold model [ahdritzOpenFoldRetrainingAlphaFold22023], extracts protein-specific embeddings $\mathcal{P}$ consisting of a single representation ($\mathbf{s}$) and a pair representation ($\mathbf{z}$), shared across frames.
+For each frame, a FrameEncoder, adapted from the template module used in prior works [jingAlphaFoldMeetsFlow2024, jumperHighlyAccurateProtein2021AF2], encodes pairwise distance of pseudo–Cβ atoms via triangular updates and merges this frame pair representations $\mathbf{z}_{\text{Frame}}^{l}$ with the protein pair representation $\mathbf{z}$.
+The resulting frame latent embedding, $\mathbf{h}^{l}=[\mathbf{s}^{l},\mathbf{z}^{l}]$, is invariant to global translation and rotation of the conformations, and are passed to the Trajectory Module.
 Following causal sequence modeling, a masked frame token “[M]” is introduced by zeroing out the pseudo-Cβ pairwise distances to remove structural information.
 Trajectory Module.
-The Trajectory Module models structural and temporal dependencies across frames, updateing each frame’s embedding based on its preceding frames.
-It consists of interleaved StructualUpdate and TemporalUpdate layers that operate on the frame-wise latent embeddings.
-StructralUpdate incorporates Pairformer layers, a core architecture in protein structure modeling, to update single and pair embeddings through triangular operations [2].
+The Trajectory Module models structural and temporal dependencies across frames, updating each frame’s embedding based on its preceding frames.
+It consists of interleaved StructuralUpdate and TemporalUpdate layers that operate on the frame-wise latent embeddings.
+StructuralUpdate incorporates Pairformer layers, a core architecture in protein structure modeling, to update single and pair embeddings through triangular operations [ahdritzOpenFoldRetrainingAlphaFold22023].
 TemporalUpdate employs a Llama-based causal transformer layer for channel-wise self-attention over the sequence of frame embeddings.
 Each channel in the single and pair embeddings is updated independently.
-Frame indices are encoded using Rotary Position Embedding [37].
+Frame indices are encoded using Rotary Position Embedding [su2023roformerenhancedtransformerrotary].
 This interleaved design enables efficient updates while maintaining flexibility in modeling sequential dependencies.
@@
 The updated latent embeddings from Trajectory Module serve as conditioning signals for generating the conformation at each frame.
-For the SE(3) diffusion model described in Section 2.3, we adopt ConfDiff [40] as the DiffusionDecoder to generate 3D conformations.
+For the SE(3) diffusion model described in Section 2.3, we adopt ConfDiff [wang2024proteinconfdiff] as the DiffusionDecoder to generate 3D conformations.
 ConfDiff composes of layers of Invariant Point Attention and Transformer (on single embeddings) to collectively update the residue SE(3) rigids, as well as single and pair embeddings.
@@
 Dataset.
-We evaluate model performance on ATLAS [39], a large-scale protein MD dataset covering $\sim$1300 proteins with diverse sizes and structures.
+We evaluate model performance on ATLAS [yannATLASProteinFlexibility2024], a large-scale protein MD dataset covering $\sim$1300 proteins with diverse sizes and structures.
 For each protein, it contains triplicated 100 ns simulation trajectories.
-All models are trained on training trajectories and evaluated on test trajectories split by protein identity [17, 40, 18].
+All models are trained on training trajectories and evaluated on test trajectories split by protein identity [jingAlphaFoldMeetsFlow2024, jing2024generative_mdgen, wang2024proteinconfdiff].
 This presents a challenging task for assessing the generalization to unseen protein structures and dynamics.
 Model training.
-All modules are trained from scratch, except for the FoldingModule (frozen OpenFold weights) and the DiffusionDecoder (initialized from ConfDiff).
-During training, sub-trajectories of length $L=8$ with varying strides ($1\sim 1024$ MD snapshots saved at 10 ps intervals) are sampled to enable learning across multiple timescales.
-For the main ConfRover model, we adopt a hybrid training strategy using a 1:1 ratio of trajectory and single-frame training objectives; For conformation interpolation task, we continue training the model with a 1:1:1 ratio of trajectory, single-frame and interpolation objectives.
+We use OpenFold (with frozen pretrained weights) as ConfRover’s FoldingModule and initialized the weights of the DiffusionDecoder from ConfDiff, while training the remaining parts of the network from scratch.
+During training, trajectories of length $L=8$ with varying timesteps (strides), corresponding to $1\sim 1024$ MD snapshots saved at 10 ps intervals, are sampled to enable learning across multiple timescales.
+For the base ConfRover, we adopt a hybrid training strategy with 1:1 ratio between trajectory and single-frame training objectives.
+To further enable conformation interpolation, we continue training the base model with a 1:1:1 ratio of trajectory, single-frame and interpolation objectives, denotes the model as ConfRover-interp.
 See Appendix D.2 for training details.
 Baselines.
-We compare ConfRover with state-of-the-art deep learning models for each task: For trajectory simulation, we compare against MDGen, a flow-based non-autoregressive trajectory model trained on ATLAS; For time-independent generation, we evaluate against AlphaFlow and ConfDiff, flow- and diffusion-based conformation generation models finetuned on ATLAS; For conformation interpolation, no existing baseline is available for large proteins, so we focus on analyzing results of our model.
-As this is an emerging research area, baseline models remain limited.
-While recent works on trajectory simulation [21, 8, 6] and conformation interpolation [10, 18] have been proposed, they are not tailored for large proteins as in ATLAS or are not publicly available at the time of manuscript preparation.
-See Appendix A for more details.
+We compare ConfRover with state-of-the-art deep learning models for each task: For trajectory simulation, we compare against MDGen [jing2024generative_mdgen], a flow-based non-autoregressive trajectory model trained on ATLAS; For time-independent generation, we evaluate against AlphaFlow [jingAlphaFoldMeetsFlow2024] and ConfDiff [wang2024proteinconfdiff], flow- and diffusion-based conformation generation models finetuned on ATLAS; For conformation interpolation, no existing baseline is available for large proteins, therefore, we focus on analyzing results of our model.
+See Appendix A for more details on the availability of baseline models.
 4.1 Trajectory Simulation
@@
 Therefore, we evaluate the model’s ability to recover trajectory dynamics from two perspectives: (1) how well it captures the magnitude of conformation changes across varying start conformations and timescales; (2) how well it recovers the conformational states and principal dynamic modes observed in long-time MD simulations.
-Table 1: Pearson correlations of conformation changes between sampled and reference trajectories in multi-start. The better scores are highlighted in bold.
 […]
-Traj. Frame $\Delta$Frame
-MDGen 0.55 0.45 0.40
-ConfRover 0.77 0.63 0.53
+Table 1: Pearson correlations between conformation changes in model-generated and reference trajectories under the multi-start setting.
+The mean and standard deviations are calculated from five independent runs.
+Models with higher correlations are highlighted in bold.
+[…]
+[…]
+[…]
+[…]
 PCA 2D
-Traj. Frame $\Delta$Frame
-MDGen 0.16 0.11 0.10
-ConfRover 0.75 0.50 0.43
+[…]
+[…]
 […]
@@
 We report the Pearson correlation of measured conformation changes between model-generated and reference trajectories in Table 1, with additional results in Appendix E.1.
-Compare with MDGen, ConfRover shows a significant improvement in correlation scores, mean absolute error and structural quality (Table 8), indicating its stronger ability to recover the magnitude of conformation changes across different starting conditions in the conformational space.
+Compared with MDGen, ConfRover shows a significant improvement in correlation scores, mean absolute error and structural quality (Table 9), indicating its stronger ability to recover the magnitude of conformation changes across different starting conditions in the conformational space.
 The greater difference observed in PCA highlights that ConfRover more accurately captures conformational changes along the feature dimensions most relevant to the structural variance observed in MD.
@@
 To confirm that this post-processing step does not introduce artifacts, we also trained MDGen models under the evaluation setups.
-The results are consistent with the downsampled version (see Appendix E.5).
+The results are consistent with the downsampled version (see Appendix E.5.1).
 Figure 4: Visualization of six proteins from multi-start.
@@
 MDGen primarily exhibits local movements, whereas ConfRover captures conformations changes similar to MD simulations.
-Table 2: State recovery in 100 ns simulation.
-Mean values over the first two PCA components are reported.
-Precision scores for all models are close to 1 and omitted.
-The better scores are highlighted in bold.
-MD 100ns is the oracle and excluded in the comparison.
+Table 2: Recovery of conformational states in the 100 ns simulation experiment.
+The mean and standard deviation are computed over five independent runs, and the better results are highlighted in bold.
+MD 100ns serves as the oracle and is excluded from the comparison.
 JSD ($\downarrow$) Recall ($\uparrow$) F1 ($\uparrow$)
 MD 100ns 0.31 0.67 0.79
-MDGen 0.56 0.30 0.44
-ConfRover 0.51 0.42 0.58
+[…]
+[…]
 Assessing long trajectory generation on 100 ns simulation.
@@
 Specifically, we discretize each principal component into 10 evenly sized “states” and measure the distribution similarity using Jensen-Shannon Distance (JSD).
-We also compute precision, recall, and F1-score on whether sample conformations fall within these known states [30, 40, 44].
+We also compute precision, recall, and F1-score on whether sample conformations fall within these known states [luStr2StrScorebasedFramework2024, wang2024proteinconfdiff, zhengPredictingEquilibriumDistributions2024aDiG].
 To evaluate dynamic mode recovery, we perform time-lagged independent component analysis (tICA) at varying lag times on both reference and sample trajectories.
@@
 (A) Correlations of principal dynamic modes between sample and reference trajectories, evaluated at varying lag time.
+The mean and standard deviation are shown as line and shadowed area, computed from five individual runs for MDGen and ConfRover.
 (B) Examples trajectories illustrating the states explored by different methods (downsampled by 5 frames for visualization).
@@
 4.2 Time-independent Conformation Sampling
-We evaluate the time-independent sampling performance of ConfRover, following the benchmark setup in Ye et al.
-[42].
-For each protein, 250 independent conformations are generated and compared to MD reference ensembles.
-Key metrics are summarized in Table 3 with full results in Appendix E.3.
-ConfRover matches the performance of state-of-the-art specialized models.
-Compared to AlphaFlow and ConfDiff, both specialized for time-independent conformation sampling, ConfRover achieves overall comparable performance and outperforms at least one of the SOTA models in five evaluation criteria.
-This demonstrate that ConfRover, despite being a general-purpose model capable of trajecotry generation, also performs strongly in sampling independent conformations that approximate the equilibrium distribution from MD simulation.
+We evaluate the time-independent sampling performance of ConfRover, following the benchmark setup in ye2024proteinbench.
+For each protein, 250 independent conformations are sampled and compared the model generated ensembles with MD reference ensembles.
+We summarize the mean and standard deviations from five independent runs with key metrics are summarized in Table 3 with full results in Appendix E.3.
+ConfRover matches the performance of state-of-the-art ensemble generation models.
+Compared with AlphaFlow and ConfDiff, ConfRover demonstrate overall comparable performance and outperforms at least one of the SOTA models in five evaluation criteria.
+This demonstrate that ConfRover, despite being a general-purpose model capable of trajectory generation, also performs strongly in sampling independent conformations that approximate the equilibrium distribution from MD simulation.
 In contrast, MDGen, which is trained solely for trajectory generation, shows suboptimal results with sequentially sampled conformations.
-Table 3: Time-independent generation benchmark.
-AlphaFlow and ConfDiff are specialized SOTA models.
-The best scores are highlighted in bold and the second-best scores are underlined.
-*MDGen does not support time-independent sampling; its sequential sampling result are evaluated.
+Table 3: Results from the time-independent generation experiment.
+AlphaFlow and ConfDiff are state-of-the-art models for direct conformation ensemble generation.
+ConfRover is the base model trained for both trajectory and time-independent generation, where as ConfRover-traj and MDGen are trained exclusively for trajectory generation.
+The mean and standard deviation are computed from five independent runs.
+The best scores are highlighted in bold, and the second-best scores are underlined.
 Pairwise
@@
 $J$ ($\uparrow$)
-AlphaFlow 0.53 0.85 2.64 1.55 2.29 0.62 0.41 0.69
-ConfDiff 0.59 0.85 2.75 1.41 2.27 0.63 0.39 0.65
-MDGen* 0.41 0.74 2.81 1.95 2.38 0.50 0.28 0.57
-ConfRover-Traj 0.47 0.85 2.83 1.41 2.30 0.53 0.36 0.60
-ConfRover 0.51 0.85 2.62 1.39 2.28 0.61 0.38 0.67
+[…]
+[…]
+[…]
+[…]
+[…]
 Effect of hybrid training.
@@
 […]
-4.4 Extension to Masked Sequence Modeling
-Figure 7: Autoregressive generation with masked language modeling.
-$\mathbf{x}^{1}$ is the initial frame for simulation.
-Masked input frames are iteratively replaced by the generated frames.
-Attention activations for the current predicting frame $\hat{\mathbf{x}}^{3}$ are highlighted in orange.
-While the causal transformers used in ConfRover enables efficient autoregressive training and generation, proposed framework can be extended to other sequence modeling paradigm, such as masked sequence modeling.
-Similar to Li et al.
-[24], sequence models trained with bidirectional attention and masked sequence modeling can perform autoregressive generation by iteratively predicting frames in a specified order.
-At each iteration, the generated frame replaces its corresponding mask and joins the input sequence for the next iteration.
-We trained a variant ConfRover-mask following this approach.
-Specifically, we replaced causal attention with bidirectional attention in the transformer and implemented a scheduled masking strategy for training, linearly increasing the mask rate from 45% to 88%.
-ConfRover-Mask learns conformation changes in trajectories but lacks inference flexibility.
-We evaluated ConfRover-mask on the multi-start benchmark.
-As shown in Table 4, ConfRover-mask outperforms the baseline, though it performs slightly worse than the ConfRover (causal) model.
-These results demonstrate the flexibility of our framework to extend to a broader range of sequence modeling approaches.
-However, masked sequence modeling presents several practical limitations.
-Similar to non-autoregressive model, ConfRover-mask lacks the flexibility to generate trajectories of variable lengths.
-We observed structural degradation when inferring with lengths differ significantly from the training setup, such as setting $L=1$ for time-independent conformation sampling.
-As a result, this approach is not well-suited for generating long trajectories without resorting to workarounds such as sliding window prediction with window sizes matching the training window size.
-While training improvements, such as varying training window sizes, may help address these issues, we leave such exploration to future work.
-Overall, the comparison between ConfRover (causal) and ConfRover-mask highlights the versatility of causal sequence modeling, which can be more readily adapted to diverse generation scenarios.
-Table 4: Results of ConfRover-mask on the Pearson correlations of conformation changes between sampled and reference trajectories in multi-start. The better scores are highlighted in bold.
+4.4 Conformation Quality
+To ensure that ConfRover generates physically plausible conformations, we further evaluate the quality using geometric assessments from MolProbity package [williams2018molprobity] and energy profiles using a coarse-grained force field MadraX [orlando2024madrax].
+We compared conformations across 38 trajectories shared between the forward simulation and interpolation experiments, including results from MDGen, ConfRover, and an oracle MD reference.
+All structures are relaxed using the refinement pipeline in OpenFold [ahdritzOpenFoldRetrainingAlphaFold22023] to enable energy comparison.
+As shown in Table 4, conformations generated by MD simulation exhibit the hightest overall quality, as expected.
+As a generative model, ConfRover also produces high-quality conformations with fewer backbone (Ramachandran) and side-chain (rotamer) outliers, more accurate covalent lengths and angles, and achieving energy levels comparable to those of MD Reference; outperforming MDGen across all metrics.
+Furthermore, we compare conformations generated from forward simulation and interpolation tasks, where the latter includes additional constraints on terminal conformations.
+ConfRover-interp shows similar geometric and energetic metrics, indicating that intermediate conformations maintains physical plausible when ConfRover tries to interpolate between two conformational states.
+Table 4: Quality of model generated conformations.
+Conformations sampled from 38 trajectories are evaluated.
+Geometric metrics are reported as mean and standard deviations across 38 cases, and energy values are reported as mean with 95% percentiles.
+The best scores are highlighted in bold.
+Ramachandran
+outliers % ($\downarrow$)
+Rotamer
+outliers % ($\downarrow$)
+Clash
+score ($\downarrow$)
+RMS
+bonds ($\downarrow$)
+RMS
+angles ($\downarrow$)
+MolProbity
+score ($\downarrow$)
+MadraX
+energy ($\downarrow$)
 […]
-Traj. Frame $\Delta$Frame
-MDGen 0.55 0.45 0.40
-ConfRover 0.77 0.63 0.53
-ConfRover-mask 0.71 0.60 0.49
-PCA 2D
-Traj. Frame $\Delta$Frame
-MDGen 0.16 0.11 0.10
-ConfRover 0.75 0.50 0.43
-ConfRover-mask 0.73 0.49 0.41
+[…]
+[…]
+[…]
 5 Conclusions and Limitations
@@
 Extensive experiments and analyses highlight several empirical advantages: (1) ConfRover outperforms the current state-of-the-art in trajectory simulation, accurately capturing dynamic magnitude, state recovery, and principal motions; (2) Despite being a multi-purpose model, it achieves competitive performance in time-independent sampling compared to specialized methods; (3) With simple sequence reordering, ConfRover effectively learns to interpolate between conformations.
-Nevertheless, several limitations still remain: (1) Trajectory simulation and interpolation are emerging tasks with limited baselines.
-We hope this work encourages further development and open-sourcing of models for comprehensive benchmarking; (2) Some evaluation metrics are still preliminary.
-While interpolation results are intuitive, they would benefit from more meaningful benchmarks reflecting functional state transitions.
-Evaluation is also constrained by the ATLAS dataset, which is limited to 100 ns simulations and mainly local dynamics.
-(3) Although ConfRover narrows the gap with classical MD, it still falls short in fully capturing the conformational space.
+Nevertheless, several limitations still remain: (1) Trajectory simulation and interpolation are emerging tasks with few available baseline models.
+We hope that this work, together with future developments in the field, will contribute to establishing more comprehensive benchmarks; (2) The dataset and evaluation metrics used in this study are limited and preliminary.
+The ATLAS dataset contains 100 ns simulations of single-chain proteins, which may not capture large conformational changes or the dynamics of protein complexes.
+Although we curated interpolation cases from ATLAS for demonstration purposes, future benchmarks reflecting realistic functional state transitions would be more meaningful.
+(3) Although ConfRover narrows the gap with classical MD, it still falls short in fully capturing the conformational space with high structural fidelity.
 Future gains may come from scaling training data, using more efficient architectures, and leveraging additional information from MD, such as energy information.
@@
 6 Acknowledgments
-We thank Zaixiang Zheng for the technical discussions and valuable feedback during the preparation of this manuscript.
-Appendix A Extended Discussions
+We would like to thank Dr.
+Hang Li for his invaluable support of this project.
+We also thank Zaixiang Zheng for insightful discussions, and Wesley Hsieh, Yi Zhou, Nima Shoghi, Yuxuan Liu, Xiaolu Shen, Jing Yuan, Yilai Li, Fei Ye, and Wei Qu for their valuable feedback.
+Appendix A Additional Background
 A.1 Related Work
 Deep Generative Models for MD Trajectories.
-Recent works have explored protein trajectory generation as a surrogate for MD simulations.
-Models such as Timewarp [21], ITO [34], and EquiJump [8] learn stochastic transport functions to sample future conformations at a lagged time (longer than MD intervals), reducing the computational costs of long-timescale simulations.
+Recent works have explored generating protein trajectories as a surrogate for MD simulations.
+Models such as Timewarp [klein2024timewarp], ITO [schreiner2024implicit], and EquiJump [costa2024equijump] learn stochastic transport functions to sample future conformations at a lagged time (longer than MD intervals), reducing the computational costs of long-timescale simulations.
 However, these methods assume Markovian dynamics by relying solely on the current state for prediction, which may not be suitable for non-Markovian dynamics common in protein MD data.
-To capture higher-order dependencies between the frames, AlphaFolding [6] incorporates history frames via “motion nodes”, but it requires a fixed context window.
-MDGen [18] takes a different approach by directly modeling the joint distributions of frames in a trajectory and learning frame dependencies through “masked frame modeling”, similar to masked language modeling.
+To capture higher-order dependencies between the frames, AlphaFolding [cheng20244d4ddiff] incorporates history frames via “motion nodes”, but it requires a fixed context window.
+MDGen [jing2024generative_mdgen] takes a different approach by directly modeling the joint distributions of frames in a trajectory and learning frame dependencies through “masked frame modeling”, similar to masked language modeling.
 However, its key-frame parameterization requires separate models for different tasks, and its non-autoregressive paradigm limits flexible generation (e.g., not compatible for generating trajectories with variable lengths).
-GST [25] applies autoregression for future frame prediction, enabling variable-length conditioning context and prediction horizons.
+GST [li2025geometricGST] applies autoregression for future frame prediction, enabling variable-length conditioning context and prediction horizons.
 While the autoregressive approach is conceptually similar to ConfRover, their work differs in several key aspects: it performs deterministic prediction rather than generative trajectory sampling; it employs a graph-based architecture with fixed structural priors from an adjacency graph, instead of full attention across all residues; it is trained and evaluated on a single protein instead of diverse proteins from ATLAS under a transferable setting.
 Beyond forward trajectory simulation, generative models have also been applied to conformation interpolation, that is, generating the intermediate trajectories between two conditioned states.
-The non-autoregressive framework of MDGen [18] can be extended to sample transition pathways between such states; however, its key-frame parameterization requires training a separate model for this task, and it has not been tested on large proteins like those in ATLAS.
-Du et al.
-[10] proposed a simulation-free objective for transition-pathway sampling based on Doob’s $h$-transform, but their approach has only been validated on numerical models and the small protein Chignolin.
+The non-autoregressive framework of MDGen [jing2024generative_mdgen] can be extended to sample transition pathways between such states; however, its key-frame parameterization requires training a separate model for this task, and it has not been tested on large proteins like those in ATLAS.
+du2024doobs proposed a simulation-free objective for transition-pathway sampling based on Doob’s $h$-transform, but their approach has only been validated on numerical models and the small protein Chignolin.
 Its generalizability to larger, more diverse proteins remains unassessed.
@@
 In contrast, ConfRover is a general framework that learns both the trajectory generation tasks as well as direct sampling of independent protein conformations.
-Deep Learning Models for Molecular Conformations.
+Deep Learning Models for Conformation Ensemble Generation.
 Another line of work focuses on direct sampling of conformations in a time-independent manner.
-Early efforts include perturbing the input to folding models (e.g., AlphaFold) [9, 41, 36] or perturbing the conformation using a structural diffusion model [30].
+Early efforts include perturbing the input to folding models (e.g., AlphaFold) [delalamoSamplingAlternativeConformational2022MSAsampling, wayment-steelePredictingMultipleConformations2024MSA, steinSPEACH_AFSamplingProtein2022MSA] or perturbing the conformation using a structural diffusion model [luStr2StrScorebasedFramework2024].
 However, these models are trained solely on static PDB structures and do not explicitly model the conformational distribution.
-Recent works have shifted to deep generative paradigms that directly learn protein-specific conformational distributions [16, 44, 17, 40, 29, 23].
+Recent works have shifted to deep generative paradigms that directly learn protein-specific conformational distributions [jingEigenFoldGenerativeProtein2023EigenFold, zhengPredictingEquilibriumDistributions2024aDiG, jingAlphaFoldMeetsFlow2024, wang2024proteinconfdiff, lu2024structureesmdiff, lewis2024scalableBioemu].
 Several models in this category, including AlphaFlow, ConfDiff, and BioEmulator, fine-tune pretrained structure models on large-scale MD datasets, enhancing their ability to capture conformational distributions.
-A related approach trains normalizing flow models to approximate the Boltzmann distribution [33, 22], but their invertibility constraints limit scalability and transferability beyond small molecules and peptides.
+A related approach trains normalizing flow models to approximate the Boltzmann distribution [noe2019boltzmann, kohler2020equivariant], but their invertibility constraints limit scalability and transferability beyond small molecules and peptides.
 While these methods can generate time-independent conformations, they overlook temporal relationships and do not capture the kinetic aspects of protein dynamics.
-ConfRover handles both time-dependent trajectory generation and time-independent conformation learning.
+Deep Learning Enhanced Molecular Dynamics.
+Another direction integrates deep learning with molecular conformation modeling through machine learning force field (MLFF) models [kovacs2025maceoff, rhodes2025orbv3atomisticsimulationscale].
+These models aim to incorporate higher-level accuracy (e.g., from ab initio calculations) into classical molecular dynamics simulations, improving fidelity while maintaining scalability.
+However, they still rely on sequential MD sampling with small integration steps and can be more computationally expensive than conventional MD.
+Two-for-One [Arts2023Twoforone] does not directly learn an MLFF but instead trains a diffusion model for protein conformations and uses the resulting score function as an approximate coarse-grained force field for MD simulation.
+Although it still depends on sequential MD sampling, this work provides an interesting perspective that connects conformation distribution modeling and trajectory generation.
+In contrast, different from these approaches, ConfRover explicitly models and generates both individual conformation and trajectories, without relying on force field-based MD simulations.
 Image-Video Generation.
 The challenge of modeling protein dynamics conceptually parallels tasks in image and video generation, requiring both data distribution learning and temporal modeling.
-Recent advances in video generation offer valuable insights into addressing these challenges.
+Recent advances in video generation offer valuable insights in addressing these challenges.
 Given limited video data, extending image generative models to video has proven effective.
-Several works [15, 5, 14] achieve this by incorporating temporal attention layers, enabling frame-to-frame communication.
+Several works [ho_video_2022, blattmann2023stablevideo, ho2022imagen] achieve this by incorporating temporal attention layers, enabling frame-to-frame communication.
 Disabling temporal attention reverts them to image models, allowing flexible training across both modalities.
@@
 Meanwhile, the extension of autoregressive language models to image and video domains has shown strong potential for sequential generation in different data modalities.
-Li et al.
-[24] integrates language models’ sequential modeling with diffusion models’ ability to model continuous distributions, showing that discrete tokens are not essential for autoregressive models.
-MarDini [27] extends the concept to video generation with efficient llama-style temporal planning and high-resolution video generation via a diffusion decoder.
+li2024autoregressivemar integrates language models’ sequential modeling with diffusion models’ ability to model continuous distributions, showing that discrete tokens are not essential for autoregressive models.
+MarDini [liu2024mardini] extends the concept to video generation with efficient llama-style temporal planning and high-resolution video generation via a diffusion decoder.
 By applying masked “frame” modeling, MarDini allows the model to learn flexible temporal relationships and enables diverse tasks such as frame interpolation.
 ConfRover differs from these works from video models in several aspects: it employs SE(3) diffusion for 3D structure generation; by using an autoregressive paradigm, it explicitly decouples the diffusion generation from temporal modeling, unlike the spatiotemporal denoising process in MarDini; in addition, the causal autoregressive framework enables more flexible trajectory generation with variable lengths.
-A.2 Training and Inference Cost
-ConfRover models in this work contains 19.6 M trainable parameters.
-All model training and sampling were carried out using 8 NVIDIA H100 GPUs with Distributed Data Parallel.
-We trained the main ConfRover model for 180 epochs ($\sim 37$ hrs) and ConfRover–interp model for additional 220 epochs ($\sim 45$ hrs).
-Inference cost varies with protein size and benchmark scale.
-For a moderately sized protein (e.g., 300 amino acids), simulating a short trajectory of 9 frames takes less than 1 minute, while generating an 80-frame trajectory (equivalent to 100 ns ATLAS setting in the paper) takes approximately 8.3 minutes.
-In comparison, a standard MD simulation under the ATLAS setup requires roughly 10 hours to reach 100 ns.
-For our large-scale benchmarks, multi-start trajectory simulation (2,700+ trajectories) took 8 hours and 30 minutes, while time-independent sampling (250 conformations per protein) took 3 hours and 20 minutes.
-A.3 Baseline Limitation
-As an emerging research area, few models currently support learning protein dynamics in transferable settings.
-Existing approaches based on forward transport operators have been primarily trained and evaluated on small peptides (e.g., Timewarp [21]) or small fast-folding proteins (e.g., EquiJump [8]).
-Although AlphaFolding [6] was trained on ATLAS, its models are not yet publicly available.
-Due to these limitations, we use MDGen as the only available model for trajectory-based comparison.
-For conformation interpolation, neither sampling-based method [10] nor video-like method MDGen [18] have been trained and evaluated on large proteins.
+A.2 Baseline Limitation
+Modeling MD trajectories using deep generative models is still an emerging research area, few models currently support learning protein dynamics in transferable settings.
+Existing approaches based on forward transport operators have been primarily trained and evaluated on small peptides (e.g., Timewarp [klein2024timewarp]) or small fast-folding proteins (e.g., EquiJump [costa2024equijump]).
+Although AlphaFolding [cheng20244d4ddiff] was also trained using ATLAS, the model weights are not publicly available at the time of this work.
+(We attempted an internal reproduction and included its result in Appendix E.5.2.) Due to these limitations, we use MDGen as the only available model for the main experiment.
+For conformation interpolation, neither sampling-based method [du2024doobs] nor video-like method MDGen [jing2024generative_mdgen] have been trained and evaluated on large proteins.
+Therefore, we focus on demonstrating the interpolation results of ConfRover-interp.
 Appendix B Diffusion Models on SE(3) Space
 Diffusion Probabilistic Models (DPM) model complex distributions through iterative denoising.
-In the context of protein conformations, DPMs defined over $\mathrm{SE(3)}$ translation-rotation space have been applied for protein backbone structural generation [43, 40].
+In the context of protein conformations, DPMs defined over $\mathrm{SE(3)}$ translation-rotation space have been applied for protein backbone structural generation [yimSEDiffusionModel2023, wang2024proteinconfdiff].
 Following Section 2.2, $\mathbf{x}_{0}=(\mathbf{T}_{0},\mathbf{R}_{0})\in\mathrm{SE(3)}^{N}$ denotes the translations and rotations of backbone rigids in data.
@@
 $\mathbf{w}_{t}$ and $\mathbf{w}_{t}^{\mathrm{SO(3)}}$ are the standard Wiener processes in $\mathcal{N}(0,I_{3})^{\otimes N}$ and $\mathcal{U}(\mathrm{SO}(3))^{\otimes N}$ respectively.
-[…]
+The transition kernel of $\mathbf{T}$ satisfies $p_{t|0}(\mathbf{T}_{t}|\mathbf{T}_{0})=\mathcal{N}(\mathbf{T}_{t};\sqrt{\alpha_{t}}\mathbf{T}_{0},(1-\alpha_{t})\mathbf{I})$, where $\alpha_{t}=e^{-\int_{0}^{t}\beta_{s}ds}$.
+The transition kernel of $\mathbf{R}$ satisfies $p_{t|0}(\mathbf{R}_{t}|\mathbf{R}_{0})=\mathcal{IGSO}_{3}(\mathbf{R}_{t};\mathbf{R}_{0},t)$, where $\mathcal{IGSO}_{3}$ is the isotropic Gaussian distribution on $\mathrm{SO(3)}$ [yimSEDiffusionModel2023].
 The associated reverse-time stochastic differential equation (SDE) follows:
@@
 These features are concatenated with the single representation from the FoldingModule.
-To encode the structural information of each conformation frame, we introduced the FrameEncoder, a pseudo-beta-carbon (Cβ) coordinate encoder similar to the InputEmbedding module from AlphaFlow [17] (without diffusion time embedding).
+To encode the structural information of each conformation frame, we introduced the FrameEncoder, a pseudo-beta-carbon (Cβ) coordinate encoder similar to the InputEmbedding module from AlphaFlow [jingAlphaFoldMeetsFlow2024] (without diffusion time embedding).
 Specifically, this module first compute the pairwise distances between residues using Cβ coordinates.
-These distances are then binned, embedded into latent embedding, and further refined through triangular update blocks including triangle attention and multiplication updates [19].
+These distances are then binned, embedded into latent embedding, and further refined through triangular update blocks including triangle attention and multiplication updates [jumperHighlyAccurateProtein2021AF2].
 See Algorithm 1 for the specifics.
@@
 Trajectory Module.
-In the Trajectory Module, we interleave layers of StructuralUpdate and TemporalUpdate to iteratively update the latent $[\mathbf{s}^{l},\mathbf{z}^{l}]$, enabling the temporal reasoning across frames and structural refinement within each frame.
-For the StructuralUpdate, we adopt a Pairformer block from AlphaFold 3 [2], which jointly updates the single and pair embeddings of the current frame through structural reasoning.
-After the StructuralUpdate, the pair embedding is flattened from $[N,N,d]$ to $[N\times N,d]$ and concatenated with the single embedding before being passed into the TemporalUpdate.
-To model temporal dependencies between frames, we use a lightweight Llama architecture [38].
+In the Trajectory Module, we interleave layers of StructuralUpdate and TemporalUpdate to iteratively update the latent $\mathbf{s}^{l},\mathbf{z}^{l}]$, enabling the temporal reasoning across frames and structural refinement within each frame.
+For the StructuralUpdate, we adopt a Pairformer block from AlphaFold 3 [ahdritzOpenFoldRetrainingAlphaFold22023], which jointly updates the single and pair embeddings of the current frame through structural reasoning.
+After the StructuralUpdate, the pair embedding is flattened from $[N,N,d]$ to $N\times N,d]$ and concatenated with the single embedding before being passed into the TemporalUpdate.
+To model temporal dependencies between frames, we use a lightweight Llama architecture [touvron2023llama].
 We transpose the input such that the temporal dimension is treated as the sequence axis for channel-wise self-attention across time.
-Rotary positional encoding [37] is applied to encode the temporal position for each frame.
+Rotary positional encoding [su2023roformerenhancedtransformerrotary] is applied to encode the temporal position for each frame.
 A causal attention mask is applied to restrict each frame to only attend to previous frames.
@@
 Trajectory Module contains interleaving blocks of StructuralUpdate and TemporalUpdate (only one block of each is shown).
-StructuralUpdate leverages the Pairformer architecture from Abramson et al.
-[1], updating the pair embeddings with triangular updates and the single embeddings using with pair bias from the updated pair.
+StructuralUpdate leverages the Pairformer architecture from abramsonAccurateStructurePrediction2024, updating the pair embeddings with triangular updates and the single embeddings using with pair bias from the updated pair.
 The updated pair embeddings are flatten and concatenate with single embedding for channel-wise temporal update.
@@
 Structure Decoder.
-Conformation generation, conditioned on the temporal signals from the Trajectory Module, is performed using the model architecture from ConfDiff [40], As shown in Figure 10, following ConfDiff, the inputs to the denoising model include diffusion time $t$, pairwise distance between residue rigids, and residue indices (not shown).
+Conformation generation, conditioned on the temporal signals from the Trajectory Module, is performed using the model architecture from ConfDiff [wang2024proteinconfdiff], As shown in Figure 9, following ConfDiff, the inputs to the denoising model include diffusion time $t$, pairwise distance between residue rigids, and residue indices (not shown).
 They are encoded and concatenated with the single and pair embeddings from FoldingModule.
@@
 D.2 Training and Inference Details
-We train all ConfRover models on the trajectories from the ATLAS training set, following the train-validation-test split of previous works [17, 40, 18].
+We train all ConfRover models on the trajectories from the ATLAS training set, following the train-validation-test split of previous works [jingAlphaFoldMeetsFlow2024, wang2024proteinconfdiff, jing2024generative_mdgen].
 Specifically, we exclude the training proteins longer than 384 amino acid residues, leading to 1080 training proteins.
@@
      Gradient Clip      1.0
-              Learning Rate               1×$\times 10^{-4}$
+[…]
      Optimizer      Adam (weight decay = 0.)
+D.3 Training and Inference Cost
+ConfRover models in this work contains 19.6 M trainable parameters.
+All model training and sampling were carried out using 8 NVIDIA H100 GPUs with Distributed Data Parallel.
+We trained the main ConfRover model for 180 epochs ($\sim 37$ hrs) and ConfRover–interp model for additional 220 epochs ($\sim 45$ hrs).
+Inference cost varies with protein size and benchmark setups.
+To measure the potential speedups from using ConfRover compared with classic MD simulation, we measured the wall-clock time required to generate 100 ns trajectories (80 frames) for ATLAS proteins of varying sizes and report the average inference time per size bucket.
+For comparison, we selected a representative protein from each bucket and estimated the time required to simulate 100 ns using OpenMM with implicit solvent.
+Both are performed on a single NVIDIA H100-80G GPU.
+As shown in the Table D.3, ConfRover provides significant speedup for 100 ns simulation, with even more pronounced acceleration for larger proteins.
+Table 8: Runtime comparison (in minutes) across different protein sequence lengths. Speedup is computed as MD runtime divided by ConfRover runtime
+Sequence length (0,150) [150,300) [300,450) [450,600) [600,724]
+ConfRover 6.99 7.53 10.92 15.88 20.83
+MD 104.54 207.92 386.69 651.29 1099.13
+[…]
+For our large-scale experimental benchmarks, multi-start trajectory simulation (2,700+ trajectories) took 8 hours and 30 minutes, while time-independent sampling (250 conformations per protein) took 3 hours and 20 minutes.
 Appendix E Additional Experimental Results
@@
 Additional Results.
-Here we also include the scatterplot of Pearson correlation in Figure 11 and additional metrics from the multi-start experiments at different strides in Table 8.
+Here we also include the scatterplot of Pearson correlation in Figure 10 and additional metrics from the multi-start experiments at different strides in Table 9 (one experimental included).
 Across different strides, ConfRover models consistently outperforms MDGen in recovering the correct level of dynamics.
-Figure 11: Scatterplots of conformation changes in sample trajectories versus those in the reference trajectories, measured by trajectory-level conformation changes, frame-level conformation changes, and next-frame difference ($\Delta$Frame), measured by the RMSD of alpha carbons (unit: Å) or $L^{2}$ distance in the projected PCA space.
-MDGen tends underestimate the magnitude of conformation changes while ConfRover generate samples at similar level as the MD reference.
-The exact match of measured conformation changes is not possible due to stochastic sampling in both MD simulation and generative models.
-Table 8: Additional metrics from the multi-start benchmark. Results for different strides are shown in separate blocks. The better score in each block is highlighted in bold (excluding diversity)
+[…]
+Table 9: Additional metrics from the multi-start benchmark.
+Results for different strides are shown in separate blocks.
+The better score in each block is highlighted in bold (excluding diversity).
+One inference run per model is used for this comparison.
 Diversity MAE on PCA-2D ($\downarrow$) MAE on CA coordinates ($\downarrow$) Quality
@@
 E.3 Time-independent Conformation Sampling
-We follow the evaluation protocol of Ye et al.
-[42] to assess time-independent conformation sampling on the ATLAS test set.
+We follow the evaluation protocol of ye2024proteinbench to assess time-independent conformation sampling on the ATLAS test set.
 For each protein, 250 independent samples are generated.
 Since MDGen does not support time-independent sampling, we approximate its performance using samples from its 100-ns trajectory, serving as a sequential-sampling baseline.
-The performance of state-of-the-art models, AlphaFlow and ConfDiff, is taken from Table 10 of Ye et al.
-[42], using their best-performing variants: AlphaFlow-MD and ConfDiff-Open-MD.
+The performance of state-of-the-art models, AlphaFlow and ConfDiff, is taken from Table 10 of ye2024proteinbench, using their best-performing variants: AlphaFlow-MD and ConfDiff-Open-MD.
 Full results are shown in Table 10.
 Table 10: Performance on time-independent conformation generation on ATLAS.
-A total of 250 conformations were sampled for each protein, and the median values across 82 proteins are reported.
-The best performance is highlighted in bold, and the second-best is underlined.*MDGen does not support time-independent sampling and the result from sequential sampling is used.
+A total of 250 conformations were sampled for each protein, and the mean and standard deviation of metrics are computed from five independent runs.
+The best performance is highlighted in bold, and the second-best is underlined.ConfRover-traj and MDGen are trained to exclusive generate trajectories.
+MDGen does not support time-independent sampling and the metrics are evaluated on sequential sampling result.
 Diversity Flexibility: Pearson $r$ on Distributional accuracy
@@
 PC sim
-$>$ 0.5 %↑
-AlphaFlow 2.87 1.63 0.53 0.66 0.85 2.64 1.55 2.29 39.0
-ConfDiff 3.43 2.21 0.59 0.67 0.85 2.75 1.41 2.27 35.4
-MDGen* 1.38 0.78 0.41 0.49 0.74 2.81 1.95 2.38 13.4
-ConfRover-traj 3.21 1.74 0.47 0.61 0.85 2.83 1.41 2.30 41.5
-ConfRover 3.63 2.23 0.51 0.64 0.85 2.62 1.39 2.28 36.6
+>> 0.5 %↑
+[…]
+[…]
+[…]
+[…]
+[…]
 Ensemble observables Quality
@@
 break % ↓
-AlphaFlow 0.62 0.41 0.69 0.35 0.0 22.2
-ConfDiff 0.63 0.39 0.65 0.33 0.5 6.5
-MDGen* 0.50 0.28 0.57 0.26 0.3 29.0
-ConfRover-traj 0.53 0.36 0.60 0.27 0.4 12.2
-ConfRover 0.61 0.38 0.67 0.32 0.5 19.2
+[…]
+[…]
+[…]
+[…]
+[…]
 E.4 Conformation Interpolation
 In conformation interpolation experiment, we selected trajectories from multi-start.
-These trajectories exhibit sufficient conformation changes (e.g., RMSD between the start and end frames $>4$) and clear interpolation path in the PCA space.
+These trajectories exhibit sufficient conformation changes (e.g., RMSD between the start and end frames >4>4 Å) and clear interpolation path in the PCA space.
 The list of selected cases are in Table 11.
@@
 Figure 15: PCA plot of 38 selected interpolation cases. MD reference trajectories and results from ConfRover-interp are shown in each plot.
-E.5 Retraining MDGen for Experiment-Specific Setups
+E.5 Extended Baseline Comparisons
+E.5.1 Retraining MDGen for Experiment-Specific Setups
 Due to its non-autoregressive design and training on fixed-length trajectories, MDGen cannot directly generate sequences of varying lengths.
@@
 To address this concern, we retrain MDGen under the evaluation settings and compare the results on the Multi-start benchmark (stride = 256) and 100 ns long trajectory simulations, as shown below in Table 12, Table 13, Table 14 and Figure 16.
-As shown in these experiments, no significant difference of performance for key metrics observed comparing MDGen with post-processed results and models specifically trained at the evaluation settings, suggesting no evident decrease of trajectory quality from the subsampling post-process.
+Experimental results show no significant difference of performance for key metrics observed comparing MDGen with post-processed results and models specifically trained at the evaluation settings, suggesting no evident decrease of trajectory quality from the subsampling post-process.
+We use one inference run per model for this experiment.
 Table 12: Compare MDGen-S256F9 with MDGen from subsampling post-process.
@@
 […]
-Trajectory Frame $\Delta$Frame
+[…]
 MDGen 0.57 0.46 0.41
@@
 PCA 2D
-Trajectory Frame $\Delta$Frame
+[…]
 MDGen 0.18 0.13 0.11
@@
 RMSD
-Traj. Frame $\Delta$Frame Traj. Frame $\Delta$Frame
+$\DeltaFrame Traj. Frame Δ\DeltaFrame$
 PepBond
@@
 MDGen-S120F80 is trained and sampled with stride of 120 MD snapshots and length of 80 frames.
+E.5.2 Internal Reproduction of AlphaFolding
+Available baseline models for protein trajectory generation are limited at the time of this work, we attempted to reproduce AlphaFolding [cheng20244d4ddiff], a diffusion-based model that generate short trajectories (i.e., blocks) and can be extended to longer through iterative generation.
+We followed the authors’ official implementation  [^2] https://github.com/fudan-generative-vision/dynamicPDB/tree/main/applications/4d_diffusion, setting the motion token count to 2 and the generation horizon (block length) to 16 frames.
+To improve generation efficiency, we increased the stride from the default 1 to 40, matching the setup in MDGen.
+Training AlphaFolding on the full ATLAS dataset resulted in out-of-memory error on NVIDIA A100-80GB GPU.
+Therefore, we adopted the authors’ filtering criterion of a maximum sequence length of 256 residues.
+Similarly, we encountered the out-of-memory error for the five largest proteins during inference.
+The model was trained for 65K steps where we saw convergence.
+For evaluation, we iteratively extended the 16-frame outputs to generate 256 frames and retained the first 250 frames to match the 100 ns simulation setting.
+We compared AlphaFolding with MDGen and ConfRover on the ATLAS 100 ns simulation task, using one inference run per model.
+As shown in Table 15, both AlphaFolding and ConfRover outperform MDGen in capturing the principal coordinates of the dominant dynamics, although AlphaFolding slightly lags behind ConfRover.
+However, as shown in Table 16, AlphaFolding produces lower-quality conformations, exhibiting inflated backbone and side-chain rotamer outliers as well as higher clash rates.
+We found more evident degradations from error accumulation (Table 17), likely due to its iterative block-wise extension, which conditions only on the last frame of the preceding block.
+In contrast, MDGen employs non-autoregressive attention across all frames, while ConfRover maintains a full attention history via KV cache.
+Owing to the noisy nature of its generated trajectories, AlphaFolding also shows inflated recall in conformational state recovery (Table 18), as the increased structural noise likely leads to artificially higher coverage in state space.
+Although our implementation is a reproduction of [cheng20244d4ddiff] and may not faithfully reflect the authors’ original model, this analysis suggests the potential limitations of block-extension-based trajectory models including AlphaFolding.
+Table 15: Pearson correlations of principal dynamic modes (PC1 and PC2) between sampled and reference trajectories, evaluated at varying lag times ($\Delta t$, in frames).
+The best scores are highlighted in bold.
+$\Delta t=1$ $\Delta t=5$ $\Delta t=10$ $\Delta t=20$
+PC 1 MDGen 0.11 0.12 0.10 0.12
+PC 1 AlphaFolding 0.15 0.15 0.16 0.18
+PC 1 ConfRover 0.19 0.17 0.19 0.19
+PC 2 MDGen 0.12 0.10 0.10 0.11
+PC 2 AlphaFolding 0.18 0.15 0.16 0.20
+PC 2 ConfRover 0.19 0.17 0.18 0.17
+Table 16: Conformation geometric quality evaluated using MolProbity. Un-relaxed conformations are used for evaluation. The best scores are highlighted in bold.
+Ramachandran
+outliers % ($\downarrow$)
+Rotamer
+outliers % ($\downarrow$)
+Clash
+score ($\downarrow$)
+RMS
+bonds ($\downarrow$)
+RMS
+angles ($\downarrow$)
+MolProbity
+score ($\downarrow$)
+AlphaFolding 2.91 15.37 151.35 0.07 4.76 3.94
+MDGen 1.87 2.85 128.08 0.04 3.14 3.28
+ConfRover 1.75 3.30 76.89 0.05 3.69 3.00
+Table 17: Average Ramachandran outliers across trajectory frame ranges.
+Compared with MDGen and ConfRover, AlphaFolding exhibits an increased level of backbone outliers when simulating longer trajectories.
+Frame range (0,15] (15,31] (31,47] (47,63] (63,79]
+AlphaFolding 1.77 2.78 3.21 3.34 3.57
+MDGen 1.59 1.85 1.96 1.95 2.05
+ConfRover 1.34 1.61 1.77 1.93 2.08
+Table 18: Recovery of conformational states in the ATLAS 100 ns simulation experiment.
+Although AlphaFolding shows the highest coverage, this may due to lower sample quality that artificially inflates diversity.
+JS-Dist $\downarrow$ Recall $\uparrow$ F1 $\uparrow$
+AlphaFolding 0.47 0.51 0.65
+MDGen 0.55 0.29 0.43
+ConfRover 0.51 0.42 0.58
+E.6 Extension to Masked Sequence Modeling
+Figure 17: Autoregressive generation with masked language modeling.
+$\mathbf{x}^{1}$ is the initial frame for simulation.
+Masked input frames are iteratively replaced by the generated frames.
+Attention activations for the current predicting frame $\hat{\mathbf{x}}^{3}$ are highlighted in orange.
+While the causal transformers used in ConfRover enables efficient autoregressive training and generation, proposed framework can be extended to other sequence modeling paradigm, such as masked sequence modeling.
+Similar to li2024autoregressivemar, sequence models trained with bidirectional attention and masked sequence modeling can perform autoregressive generation by iteratively predicting frames in a specified order.
+At each iteration, the generated frame replaces its corresponding mask and joins the input sequence for the next iteration.
+We trained a variant ConfRover-mask following this approach.
+Specifically, we replaced causal attention with bidirectional attention in the transformer and implemented a scheduled masking strategy for training, linearly increasing the mask rate from 45% to 88%.
+ConfRover-Mask learns conformation changes in trajectories but lacks inference flexibility.
+We evaluated ConfRover-mask on the multi-start benchmark.
+As shown in Table 19, ConfRover-mask outperforms the baseline, though it performs slightly worse than the ConfRover (causal) model.
+These results demonstrate the flexibility of our framework to extend to a broader range of sequence modeling approaches.
+However, masked sequence modeling presents several practical limitations.
+Similar to non-autoregressive model, ConfRover-mask lacks the flexibility to generate trajectories of variable lengths.
+We observed structural degradation when inferring with lengths differ significantly from the training setup, such as setting $L=1$ for time-independent conformation sampling.
+As a result, this approach is not well-suited for generating long trajectories without resorting to workarounds such as sliding window prediction with window sizes matching the training window size.
+While training improvements, such as varying training window sizes, may help address these issues, we leave such exploration to future work.
+Overall, the comparison between ConfRover (causal) and ConfRover-mask highlights the versatility of causal sequence modeling, which can be more readily adapted to diverse generation scenarios.
+Table 19: Results of ConfRover-mask on the Pearson correlations of conformation changes between sampled and reference trajectories in multi-start. The better scores are highlighted in bold.
+[…]
+[…]
+MDGen 0.55 0.45 0.40
+ConfRover 0.77 0.63 0.53
+ConfRover-mask 0.71 0.60 0.49
+PCA 2D
+[…]
+MDGen 0.16 0.11 0.10
+ConfRover 0.75 0.50 0.43
+ConfRover-mask 0.73 0.49 0.41